### v1

Abstract
We study the common continual learning setup where an overparameterized model is sequentially fitted to a set of jointly realizable tasks. We analyze the forgetting—loss on previously seen tasks—after $k$ iterations. For linear models, we prove that fitting a task is equivalent to a single stochastic gradient descent (SGD) step on a modified objective. We develop novel last-iterate SGD upper bounds in the realizable least squares setup, and apply them to derive new results for continual learning. Focusing on random orderings over $T$ tasks, we establish universal forgetting rates, whereas existing rates depend on the problem dimensionality or complexity. Specifically, in continual regression with replacement, we improve the best existing rate from $\mathcal{O}((d-\bar{r})/k)$ to $\mathcal{O}(\min(1/\sqrt[4]{k},\sqrt{d-\bar{r}}/k,\sqrt{T\bar{r}}/k))$, where $d$ is the dimensionality and $\bar{r}$ the average task rank. Furthermore, we establish the first rates for random task orderings without replacement. The obtained rate of $\mathcal{O}(\min(1/\sqrt[4]{T},\,(d-r)/T))$ proves for the first time that randomization alone—with no task repetition—can prevent catastrophic forgetting in sufficiently long task sequences. Finally, we prove a similar $\mathcal{O}(1/\sqrt[4]{k})$ universal rate for the forgetting in continual linear classification on separable data. Our universal rates apply for broader projection methods, such as block Kaczmarz and POCS, illuminating their loss convergence under i.i.d. and one-pass orderings.
1 Introduction
In continual learning, tasks are presented sequentially, one at a time. The goal is for the learner to adapt to the current task—e.g., by fine-tuning using gradient-based algorithms—while retaining knowledge from previous tasks. A central challenge in this setting is termed catastrophic forgetting, where expertise from earlier tasks is lost when adapting to newer ones. Forgetting is influenced by factors such as task similarity and overparameterization (Goldfarb et al., 2024), and is also related to trade-offs like the plasticity-stability dilemma (Mermillod et al., 2013). Continual learning is becoming increasingly important with the rise of foundation models, where retraining is prohibitively expensive and data from prior tasks is often unavailable, e.g., due to privacy concerns or data retention issues.
Previous work has shown, both analytically (Evron et al., 2022, 2023; Kong et al., 2023; Jung et al., 2025; Cai and Diakonikolas, 2025) and empirically (Lesort et al., 2023; Hemati et al., 2024), that forgetting diminishes over time when task ordering is cyclic or random. Different orderings can be explored from multiple perspectives: as a strategy to mitigate forgetting (e.g., by actively ordering an agent’s learning environments); as a naturally occurring phenomenon, such as periodic trends in e-commerce systems; or as a means to model and analyze widely used continual learning benchmarks, such as randomly split datasets.
Our work focuses on a widely studied analytical setting—realizable continual linear regression, where $T$ tasks are learned sequentially over $k$ iterations according to a uniform random ordering. Evron et al. (2022) established that the worst-case expected forgetting lies between $\Omega\left({{1}/{k}}\right)$ and $\mathcal{O}\left({{(d-\bar{r})}/{k}}\right)$, where $d$ is the problem dimensionality, and $\bar{r}$ is the average rank of individual data matrices. This raises a fundamental question with critical implications in highly overparameterized regimes: Does the worst-case forgetting necessarily scale with the dimensionality, and if so, is the dependence indeed linear?
To this end, we bridge continual learning and the literature on last-iterate stochastic gradient descent (SGD) analysis. We revisit an established connection between continual linear regression and the Kaczmarz method for solving systems of linear equations (Kaczmarz, 1937; Evron et al., 2022). Given rank-$1$ tasks, this method is known to perform a normalized stochastic gradient step on the least squares objective, fully minimizing the current task’s loss and implying a “stepwise-optimal” step size. Deepening this connection, we prove that even for general data ranks, learning a task in continual linear regression and performing an update in the Kaczmarz method, are both equivalent to a single SGD step on a modified objective with a constant, stepwise-optimal step size.
Motivated by this, we prove convergence rates for the last iterate of fixed-step size SGD that, crucially, hold for a broad range of step sizes not covered by prior work (e.g., Shamir and Zhang, 2013; Ge et al., 2019; Berthier et al., 2020; Zou et al., 2021; Varre et al., 2021; Wu et al., 2022). Somewhat curiously, prior results either hold only for the average iterate (e.g., Bach and Moulines, 2013), or for small step sizes, bounded away from the stepwise-optimal step size crucial for the continual learning setup (e.g., Varre et al., 2021). We overcome this challenge through a careful combination of analysis techniques for SGD (Srebro et al., 2010; Shamir and Zhang, 2013), further tightening the analysis to accommodate a wider range of step sizes, including the stepwise-optimal one.
Applying our last-iterate analysis to continual regression, we tighten the existing forgetting rate and also establish the first dimension-independent rate (see Table 1). Further, we provide the first rate for random task orderings without replacement, proving that task repetition is not required to guarantee convergence when ${k=T\to\infty}$, thus highlighting the effect of randomization as compared to repetition of tasks. Our results also yield novel rates for the closely related Kaczmarz and Normalized Least Mean Squares methods.
Finally, we prove a matching rate for the squared loss of the broader Projection Onto Convex Sets framework (POCS; see Gubin et al., 1967). This extends our results to continual linear classification on separable data—previously linked to projection algorithms (Evron et al., 2023)—and provides this setting’s first universal rate, independent of the problem’s “complexity”.
Summary of Contributions.
To summarize, our main contributions in this paper are as follows:
  • We establish a new reduction from continual linear regression to SGD with a particular choice of a “stepwise-optimal” step size, generalizing ideas from recent work that only applied to rank-$1$ tasks, to tasks of arbitrary rank. This facilitates last iterate-analysis to study forgetting.
  • We provide novel last-iterate convergence analysis for SGD in a realizable least squares setup. To the best of our knowledge, this is the first analysis that provides nontrivial rates for large step sizes, which are crucial to the reduction to continual learning scenarios.
  • Our main contribution, building on these techniques, is a set of improved rates of forgetting in continual linear regression, including the first universal rates, independent of the problem dimensionality or complexity, as well as the first rates for without-replacement orderings, indicating task repetition is not mandatory to diminish forgetting. See Table 1 for more details.
  • We further relate and extend our results to other settings, including continual linear classification, the block Kaczmarz method, and the Projection Onto Convex Sets framework (POCS).
2 Main Setting: Continual Linear Regression
We mainly investigate the fundamental continual linear regression setting, as studied in many theoretical papers. This setting is easy-to-analyze, yet often sheds light on important phenomena in continual learning (e.g., Doan et al., 2021; Evron et al., 2022; Lin et al., 2023; Peng et al., 2023; Goldfarb and Hand, 2023; Li et al., 2023; Goldfarb et al., 2024; Hiratani, 2024).
Notation.
[…]
Formally, the learner has access to a task collection of $T$ finite-dimensional linear regression tasks, i.e., $(\mathbf{X}_{1},\mathbf{y}_{1}),\dots,(\mathbf{X}_{T},\mathbf{y}_{T})$ where $\mathbf{X}_{m}\in\mathbb{R}^{n_{m}\times d},\mathbf{y}_{m}\in\mathbb{R}^{n_{m}}$. For $k$ iterations, the learner learns the tasks according to a task ordering $\tau:\left[{k}\right]\to\left[{T}\right]$. We analyze random orderings, previously studied in continual linear models (e.g., Evron et al., 2022, 2023; Jung et al., 2025).
Definition 1 (Random Task Ordering).
A random ordering selects tasks uniformly at random from the task collection $\left[{T}\right]$, i.e., $\tau(1),\dots,\tau(k)\sim\textrm{Unif}\left({\left[{T}\right]}\right)$, with or without replacement.
We are now ready to define the learning scheme we study, which, at each iteration, naively minimizes the sum of squared errors for the current regression task.111 This objective is natural for regression tasks; our analysis also extends to the mean squared error (by adjusting $R$ in the results).
Scheme 1 Continual Linear Regression (to Convergence)
  Initialize $\mathbf{w}_{0}=\mathbf{0}_{d}$
   For each task $t=1,\dots,k$:
       $\mathbf{w}_{t}\leftarrow$ Start from $\mathbf{w}_{t-1}$ and minimize the current loss $\mathcal{L}_{\tau(t)}(\mathbf{w})\triangleq{\frac{1}{2}}\left\|{\mathbf{X}_{\tau({t})}\mathbf{w}-\mathbf{y}_{\tau({t})}}\right\|^{2}$           with (S)GD to convergence [^2] Learning to convergence facilitates the analysis, but other analytical choices exist (see Jung et al., 2025).
   Output $\mathbf{w}_{k}$
This scheme was previously linked to the Kaczmarz method and, in a special case, to normalized SGD (Evron et al., 2022). In Section 3, we explain and develop these connections to enable novel analysis.
Our primary assumption is that there exist offline solutions that perfectly solve all $T$ tasks jointly, as assumed in much of the theoretical continual learning literature (e.g., Evron et al., 2022, 2023; Goldfarb et al., 2024; Jung et al., 2025). This assumption simplifies the analysis and rules out cases where forgetting previous tasks is beneficial, as new tasks may directly contradict them. Finally, this assumption is reasonable in highly overparameterized models and is thus linked to the linear dynamics of deep neural networks in the neural tangent kernel (NTK) regime (e.g., see Jacot et al., 2018).
Assumption 1 (Joint Linear Realizability).
We assume the set of offline solutions that solve all tasks is nonempty. That is, $\mathcal{W}_{\star}\triangleq\Big{\{}\mathbf{w}\in\mathbb{R}^{d}~{}\Big{|}~{}\mathbf{X}_{m}\mathbf{w}=\mathbf{y}_{m},~{}\forall m\in\left[{T}\right]\Big{\}}\neq\varnothing\,.$
To facilitate the results and discussions in our paper, we focus on the offline solution with minimal norm, often associated with good generalization capabilities.
Definition 2 (Minimum-Norm Offline Solution).
We specifically denote,
$\mathbf{w}_{\star}\triangleq{\operatorname*{argmin}}_{\mathbf{w}\in\mathcal{W}_{\star}}\left\|{\mathbf{w}}\right\|\,.$
Commonly in continual learning setups, the model performance on past tasks degrades, sometimes significantly, even in linear models (Evron et al., 2022). Our goal is to bound the worst-case degradation, i.e., “forgetting”, in linear models under random task orderings. We follow common definitions (e.g., Doan et al., 2021; Evron et al., 2023), and define forgetting as the average increase in the loss of the last iterate on previous tasks. In realizable settings, forgetting takes the form below.
Definition 3 (Forgetting).
Let $\mathbf{w}_{1},\dots,\mathbf{w}_{k}$ be the iterates of Scheme 1 under a task ordering $\tau$. The forgetting at iteration $k$ is the average increase in the loss of previously seen tasks. In our realizable setting, the forgetting becomes an in-sample loss. Formally,
$F_{\tau}(k)=\frac{1}{k}\sum_{t=1}^{k}\big{(}{\mathcal{L}_{\tau(t)}(\mathbf{w}_{k})-\underbrace{\mathcal{L}_{\tau(t)}(\mathbf{w}_{t})}_{=0}}\big{)}=\frac{1}{2k}\sum_{t=1}^{k}\left\|{\mathbf{X}_{\tau(t)}\mathbf{w}_{k}-\mathbf{y}_{\tau(t)}}\right\|^{2}\,.$
Under arbitrary orderings, Evron et al. (2022) showed forgetting can be “catastrophic” in the sense that $\lim_{k\to\infty}\mathbb{E}\left[F_{\tau}(k)\right]>0$. However, as we show, this cannot happen under the random ordering.
Remark 4 (Forgetting vs. Regret).
While regret and forgetting are related, they can differ significantly (Evron et al., 2022). Regret is a key quantity in online learning, defined as $\frac{1}{2k}\sum_{t=1}^{k}\|{\mathbf{X}_{\tau(t)}\mathbf{w}_{t-1}-\mathbf{y}_{\tau(t)}}\|^{2}$ in our setting. That is, it measures the suboptimality of each iterate on the consecutive task. In contrast, forgetting evaluates an iterate’s performance across all earlier tasks.
We further define the training loss to easily discuss links to other fields, such as Kaczmarz.
Definition 5 (Training Loss).
The training loss of any vector $\mathbf{w}\in\mathbb{R}^{d}$ is given by,
$\mathcal{L}(\mathbf{w})$ $=\frac{1}{T}\sum_{m=1}^{T}\mathcal{L}_{m}(\mathbf{w})=\frac{1}{2T}\sum_{m=1}^{T}\left\|{\mathbf{X}_{m}\mathbf{w}-\mathbf{y}_{m}}\right\|^{2}\,.$
Our results bound both the expected forgetting and expected training loss, leveraging a key property: expected (in-sample) forgetting can be upper bounded using expected training loss across all tasks. Specifically, Lemma 16 (in Appendix B) states that $\mathbb{E}_{\tau}[F_{\tau}(k)]\leq 2\mathbb{E}_{\tau}\left[\mathcal{L}\left(\mathbf{w}_{k-1}\right)\right]+\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}$ in orderings with replacement, where $R\triangleq\max_{m\in\left[{T}\right]}\left\|{\mathbf{X}_{m}}\right\|$ is the data “radius” and the dependence of $\mathbf{w}_{k-1}$ on $\tau_{1},\dots,\tau_{k-1}$ is implicit. Without-replacement orderings yield a related but more refined bound. The additive $\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}$ term is negligible compared to the bounds obtained in our work.
3 Reductions: From Continual Linear Regression to Kaczmarz to SGD
Previous work established connections between continual linear regression and the Kaczmarz method in the realizable case (Evron et al., 2022). We revisit these connections pedagogically to ensure our paper is self-contained. Importantly, this leads to novel links between continual learning, the Kaczmarz method, and SGD on special functions (Schemes 1,2,3), allowing us to improve the rates for continual and Kaczmarz methods by analyzing the last iterate of SGD algorithms instead.
Scheme 2 The Block Kaczmarz Method
  Input: Jointly realizable ${(\mathbf{X}_{m},\mathbf{y}_{m}),\forall m\in\left[{T}\right]}$
  Initialize $\mathbf{w}_{0}=\mathbf{0}_{d}$
   For each iteration $t=1,\dots,k$:
       $\mathbf{w}_{t}\leftarrow\mathbf{w}_{t-1}-\mathbf{X}_{\tau(t)}^{+}({\mathbf{X}_{\tau(t)}\mathbf{w}_{t-1}-\mathbf{y}_{\tau(t)}})$
Scheme 3 SGD with $\eta=1$ on special $\{f_{m}\}_{m}$
  Input: ${f_{m}(\mathbf{w})=\frac{1}{2}\left\|{\mathbf{X}_{m}^{+}\mathbf{X}_{m}\left({\mathbf{w}-\mathbf{w}_{\star}}\right)}\right\|^{2},\forall m\in\left[{T}\right]}$
  Initialize $\mathbf{w}_{0}=\mathbf{0}_{d}$
   For each iteration $t=1,\dots,k$:
       $\mathbf{w}_{t}\leftarrow\mathbf{w}_{t-1}-\nabla_{\mathbf{w}}f_{\tau(t)}\big{(}{\mathbf{w}_{t-1}}\big{)}$
3.1 Revisit: Continual Linear Regression and the Kaczmarz Method
The (block) Kaczmarz method in Scheme 2 (Kaczmarz, 1937; Elfving, 1980) is a classical iterative method for solving a linear system $\mathbf{X}\mathbf{w}=\mathbf{y}$, easily mapped to our learning problem by stacking tasks in blocks, i.e.,
$\mathbf{X}=\begin{pmatrix}\mathbf{X}_{1}\\ \vdots\\ \mathbf{X}_{T}\end{pmatrix}\in\mathbb{R}^{N\times d},\quad\mathbf{y}=\begin{pmatrix}\mathbf{y}_{1}\\ \vdots\\ \mathbf{y}_{T}\end{pmatrix}\in\mathbb{R}^{N},\quad\text{where $N=\sum_{m=1}^{T}n_{m}$.}$
In each iteration, the Kaczmarz method (Scheme 2) perfectly solves the current block, i.e., ${\mathbf{X}_{\tau(t)}\mathbf{w}_{\tau(t)}=\mathbf{y}_{\tau(t)}}$ (to see that, recall that $\mathbf{X}_{\tau(t)}^{+}$ denotes the Moore-Penrose pseudo-inverse of $\mathbf{X}_{\tau(t)}^{+}$).
The continual Scheme 1 also minimizes the current loss to convergence, i.e., until it is perfectly solved (in the realizable case). In fact, Evron et al. (2022) identified the following reduction.
Reduction 1 (Continual Regression $\Rightarrow$ Block Kaczmarz).
In the realizable case (Assumption 1) under any ordering $\tau$, continual linear regression learned to convergence333 The continual learner minimizes $\mathcal{L}_{\tau(t)}$ with (S)GD to convergence; they do not explicitly compute pseudo-inverses. is equivalent to the block Kaczmarz method. That is, the iterates $\mathbf{w}_{0},\dots,\mathbf{w}_{k}$ of Schemes 1 and 2 coincide.
3.2 New Reduction: Kaczmarz Method and Stepwise-Optimal Stochastic Gradient Descent
Rank-$1$ data.
It is known that when each task contains just one row, each update in the Kaczmarz method corresponds to a gradient step on with a specific “normalizing” step size (Needell et al., 2014). That is, noticing that in rank-$1$ we have $\mathcal{L}_{\tau(t)}(\mathbf{w})=\frac{1}{2}\big{\|}{\mathbf{x}_{\tau({t})}^{\top}\mathbf{w}-y_{\tau({t})}}\big{\|}^{2}$, the Kaczmarz updates hold
$\mathbf{w}_{t}=\mathbf{w}_{t-1}-\frac{1}{\|{\mathbf{x}_{\tau(t)}}\|^{2}}\big{(}{\mathbf{x}_{\tau(t)}^{\top}\mathbf{w}_{t-1}-y_{\tau(t)}}\big{)}\mathbf{x}_{\tau(t)}=\mathbf{w}_{t-1}-\frac{1}{\|{\mathbf{x}_{\tau(t)}}\|^{2}}\nabla_{\mathbf{w}}\mathcal{L}_{\tau(t)}(\mathbf{w}_{t-1})\,.$ (1)
What about higher data ranks?
We now establish a more general reduction from the block Kaczmarz method—at any rank—to SGD (in Section 6, we similarly connect SGD and the broader Projection Onto Convex Sets framework, extending our results to continual linear classification).
Reduction 2 (Block Kaczmarz $\Rightarrow$ SGD).
In the realizable case (Assumption 1) under any ordering $\tau$, the block Kaczmarz method is equivalent to SGD with a step size of $\eta=1$, applied w.r.t. a convex, $1$-smooth least squares objective: $\big{\{}f_{m}(\mathbf{w})\triangleq\frac{1}{2}\left\|{\mathbf{X}_{m}^{+}\mathbf{X}_{m}\left({\mathbf{w}-\mathbf{w}_{\star}}\right)}\right\|^{2}\big{\}}_{m=1}^{T}$. That is, the iterates $\mathbf{w}_{0},\dots,\mathbf{w}_{k}$ of Schemes 2 and 3 coincide.
The reduction follows from the next lemma, which states the convexity and smoothness of $f_{m}$ and expresses the gradient $\nabla_{\mathbf{w}}f_{\tau(t)}$, subsequently substituted into $\big{(}{\mathbf{w}_{t-1}-\nabla_{\mathbf{w}}f_{\tau(t)}(\mathbf{w}_{t-1})}\big{)}$ to complete the proof.
Lemma 6 (Properties of the Modified Objective).
Consider any realizable task collection such that ${\mathbf{X}_{m}\mathbf{w}_{\star}=\mathbf{y}_{m}},\forall m\in\left[{T}\right]$. Define $f_{m}(\mathbf{w})=\frac{1}{2}\left\|{\mathbf{X}_{m}^{+}\mathbf{X}_{m}\left({\mathbf{w}-\mathbf{w}_{\star}}\right)}\right\|^{2}$. Then, $\forall m\in\left[{T}\right],\mathbf{w}\in\mathbb{R}^{d}$
  1. (i)
    Upper bound: $\mathcal{L}_{m}(\mathbf{w})\leq R^{2}f_{m}(\mathbf{w})\triangleq\max_{m^{\prime}\in\left[{T}\right]}\left\|{\mathbf{X}_{m^{\prime}}}\right\|^{2}f_{m}$
  2. (ii)
    Gradient:       $\nabla_{\mathbf{w}}f_{m}(\mathbf{w})=\mathbf{X}_{m}^{+}\mathbf{X}_{m}\left({\mathbf{w}-\mathbf{w}_{\star}}\right)=\mathbf{X}_{m}^{+}\mathbf{X}_{m}\mathbf{w}-\mathbf{X}_{m}^{+}\mathbf{y}_{m}$
  3. (iii)
    Convexity and Smoothness: $f_{m}$ is convex and $1$-smooth.
The proof is given in Appendix B.
4 Rates for Random-Order Continual Linear Regression and Kaczmarz
This section focuses on improving the best upper bound known in prior continual learning literature on random orderings, summarized in Table 1. Specifically, for with-replacement random orderings, Evron et al. (2022) proved a forgetting rate of $\mathbb{E}_{\tau}\left[F_{\tau}(k)\right]=\mathcal{O}\!\left({\frac{d-\bar{r}}{k}}\right)$ where $\bar{r}\triangleq\frac{1}{T}\sum_{m}\operatorname{rank}(\mathbf{X}_{m})$. Notably, this rate depends on the problem dimensionality $d$, raising concerns when generalizing insights from linear models to deep neural networks—which are often highly overparameterized (e.g., in the NTK regime). Encouragingly, that paper only provided (implicitly) a $1/k$ lower bound for the worst-case forgetting, calling for further research into narrowing this gap.
We tighten the existing rate’s problem-dependent term from $\big{(}{d-\bar{r}}\big{)}$ to $\min\big{(}{\sqrt{d-\bar{r}},\sqrt{T\bar{r}}}\big{)}$ and also prove a problem-independent rate of ${1/\sqrt[4]{k}}$. Finally, we provide the first rates for without-replacement orderings, isolating the effect of randomness versus repetition.
Table 1: Forgetting and Loss Rates in Continual Linear Regression (and Block Kaczmarz). Upper bounds apply to any $T$ realizable tasks (or blocks). Lower bounds indicate worst cases, i.e., specific constructions. Random ordering bounds apply to the expected forgetting (or loss).
We omit mild constant multiplicative factors and an unavoidable $\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}$ term. Finally, ${a\!\wedge\!b\triangleq\min(a,b)}$. Recall: $k=\,$iterations; $d=\,$dimensionality; $\bar{r},r_{\max}=\,$average and maximum data matrix ranks.
Paper / Ordering Bound Random with Replacement Random w/o Replacement Cyclic
Evron et al. (2022) Upper $\frac{d-\bar{r}}{k}$ $\frac{T^{2}}{\sqrt{k}}\!\wedge\!\frac{T^{2}(d\!-\!r_{\max})}{k}$
Kong et al. (2023) Upper $\frac{T^{3}}{k}$
Ours (2025) Upper $\frac{1}{\sqrt[4]{k}}\wedge\frac{\sqrt{d-{\bar{r}}}}{k}\wedge\frac{\sqrt{T\bar{r}}}{k}$ $\frac{1}{\sqrt[4]{T}}\wedge\frac{d-{\bar{r}}}{T}$
Evron et al. (2022) Lower $\frac{1}{k}$ (*) $\frac{1}{T}$ (*) $\frac{T^{2}}{k}$
(*) They did not explicitly provide such lower bounds, but the $T\!=\!2$ tasks construction from their proof of Theorem 10, can yield a $\Theta({1}/{k})$ random behavior by cloning those $2$ tasks $\left\lfloor{T/2}\right\rfloor$ times for any general $T$.
4.1 A Parameter-Dependent $\mathcal{O}(1/k)$ Rate
Here, we present a tighter $\sqrt{d-\bar{r}}$ term and a term depending only on the rank and number of tasks.
Theorem 7 (Parameter-Dependent Forgetting Rate for Random With Replacement).
Under a random ordering with replacement over $T$ jointly realizable tasks, the expected loss and forgetting of Schemes 12 after $k\geq 3$ iterations are bounded as,
$\mathbb{E}_{\tau}\left[\mathcal{L}\left(\mathbf{w}_{k}\right)\right]$ $\leq\frac{\min\left({\sqrt{d-\bar{r}},\sqrt{T\bar{r}}}\right)\|{\mathbf{w}_{\star}}\|^{2}R^{2}}{2e(k-1)}\,,\quad\quad\mathbb{E}_{\tau}\left[F_{\tau}(k)\right]\leq\frac{3\min\left({\sqrt{d-\bar{r}},\sqrt{T\bar{r}}}\right)\|{\mathbf{w}_{\star}}\|^{2}R^{2}}{2\left({k-2}\right)}\,,$
where $\bar{r}\triangleq\frac{1}{T}\sum_{m\in[T]}\operatorname{rank}(\mathbf{X}_{m})$. (Recall that ${R\triangleq\max_{m\in[T]}\left\|{\mathbf{X}_{m}}\right\|}$.)
The $\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}$ factor is a scaling term, generally unavoidable in this area. The proof is outlined below, and detailed in Appendix C. Our analysis is related to a recent work (Guo et al., 2022) that also analyzes a linear map to characterize the weak error (resembling our loss). They achieve polynomial rates, that, unlike ours, involve matrix properties related to the condition number.
Proof Idea.
We rewrite the Kaczmarz update (Scheme 2) in a recursive form of the differences, i.e., $\mathbf{w}_{t}-\mathbf{w}_{\star}={\mathbf{P}}_{\tau(t)}\left({\mathbf{w}_{t-1}-\mathbf{w}_{\star}}\right)$, with a suitable projection matrix ${\mathbf{P}}_{\tau(t)}$. We define the linear map $Q\left[{\mathbf{A}}\right]=\frac{1}{T}\sum_{m=1}^{T}{\mathbf{P}}_{m}\mathbf{A}{\mathbf{P}}_{m}$ to capture the evolution of the difference’s second moments, enabling sharp analysis of the expected loss in terms of $Q$. Using properties of $Q$, norm inequalities, and the spectral mapping theorem, we establish a fast $\mathcal{O}\left({1/k}\right)$ rate with explicit dependence on $T$, $d$, and $\bar{r}$.
4.2 A Universal $\mathcal{O}(1/\sqrt[4]{k})$ Rate
Next, we present a forgetting rate independent on the dimensionality, rank, and number of tasks. This is crucial in highly overparameterized regimes which are connected to deep neural networks.
Theorem 8 (Universal Forgetting Rate for With-Replacement Random Ordering).
Under a random ordering with replacement over $T$ jointly realizable tasks, the expected loss and forgetting of Schemes 12 after $k\geq 2$ iterations are bounded as,
[…]
We prove this result in Section D.1, by leveraging the connections between continual learning and SGD. Specifically, in Section 3 we showed that continual linear regression and the Kaczmarz method, are equivalent to SGD with a step size of exactly $1$ on a related least squares problem. Our result follows from our novel last-iterate SGD bounds that, crucially, apply even to that specific step size. To ease readability, here we focused on a continual learning perspective, deferring last-iterate analysis to Section 5.
4.3 Random Task Orderings Without Replacement
Evron et al. (2022) suggested defining forgetting as “catastrophic” only when $\lim_{k\to\infty}\mathbb{E}\left[F_{\tau}(k)\right]>0$.They presented such an adversarial case with a deterministic task ordering where $k=T\to\infty$. They showed that task recurrence, under cyclic or random orderings, mitigates forgetting. So far, in random orderings, it was hard to isolate the effect of randomness from that of repetitions. It was thus unclear whether catastrophic forgetting can be alleviated by randomly permuting the tasks. Below, we provide the first result demonstrating that no recurrence is needed under random orderings.
Theorem 9 (Forgetting Rates for Without-Replacement Random Ordering).
Under a random ordering without replacement over $T$ jointly realizable tasks, the expected loss and forgetting of Schemes 12 after $k\in\{2,\dots,T\}$ iterations are both bounded as,
$\mathop{\mathbb{E}}_{\tau}\left[\mathcal{L}\left(\mathbf{w}_{k}\right)\right],~{}\mathop{\mathbb{E}}_{\tau}\left[F_{\tau}(k)\right]\leq\min\left(\frac{7}{\sqrt[4]{k-1}},\,\frac{d-\bar{r}+1}{k-1}\right){\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}.$
The proof of the dimensionality-dependent term is similar to the one of the with-replacement case, given in Section D.1.2 of Evron et al. (2022), but requires a more careful upper bound on the (in-sample) forgetting. The proof of the dimensionality-independent term again relies on last-iterate analysis, as presented in Section 5.2. Both proofs are given in Section D.2.
In Appendix A, we discuss relations between without-replacement random orderings and other areas like shuffle SGD.
5 Last-Iterate SGD Bounds for Linear Regression
In this self-contained section, we derive last-iterate guarantees for SGD in the realizable stochastic least squares setup. Motivated by the connection with continual regression discussed in Section 3, we focus on regression problems that are $\beta$-smooth individually, and obtain upper bounds for the last SGD iterate that apply for a significantly wider range of step sizes compared to prior art (Varre et al., 2021). Notably, this is the first time convergence of SGD in this setup is established for a range of step sizes completely independent of the optimization horizon. Table 2 compares our bounds with related work and classic results.
Many recent works study SGD in the realizable least squares setting (Ge et al., 2019; Vaswani et al., 2019; Berthier et al., 2020; Zou et al., 2021; Varre et al., 2021; Wu et al., 2022) under a slightly more general noise model. These studies are primarily motivated by the connection between this setup and deep neural networks in the overparameterized regime (Ma et al., 2018), where the model is expressive enough to perfectly fit the training data. With the exception of Varre et al. (2021), most of these works focus on non-fixed step size schedules and/or provide guarantees for the average iterate. Further discussion of related work can be found in Appendix A. Proceeding, throughout this section we consider stochastic, realizable least squares problem as defined next.
Setup 1.
Let $\mathcal{I}$ be an index set, and $\mathcal{D}$ a distribution over $\mathcal{I}$. We consider the optimization objective:
${\min}_{\mathbf{w}\in\mathbb{R}^{d}}\left\{\,\bar{f}(\mathbf{w})\triangleq\mathbb{E}_{i\sim\mathcal{D}}f(\mathbf{w};i)\triangleq\mathbb{E}_{i\sim\mathcal{D}}\left[\tfrac{1}{2}\left\|{\mathbf{A}_{i}\mathbf{w}-\mathbf{b}_{i}}\right\|^{2}\right]\,\right\},$
where $\mathbf{A}_{i}\in\mathbb{R}^{n_{i}\times d},\mathbf{b}_{i}\in\mathbb{R}^{n_{i}},~{}\forall i\in\mathcal{I}$. We specifically focus on $\beta$-smooth functions, i.e., $\left\|{\mathbf{A}_{i}}\right\|\leq\beta,\forall i\in\mathcal{I}$, under a realizable assumption, i.e., $\exists\mathbf{w}_{\star}\in\mathbb{R}^{d}:\bar{f}(\mathbf{w}_{\star})=0$.
Table 2: State-of-the-art Loss Bounds for Fixed-Step-Size SGD. We consider stochastic convex optimization with an objective $\bar{f}(\mathbf{w})\!\triangleq\!\mathbb{E}_{\xi}f(\mathbf{w};\xi)$, where $f(\cdot;\xi)$ is $\beta$-smooth almost surely, $\sigma^{2}\geq\mathbb{E}\|\nabla f(\mathbf{w};\xi)-\nabla\bar{f}(\mathbf{w})\|^{2}$, $\sigma_{\star}^{2}\triangleq\mathbb{E}\|\nabla f(\mathbf{w}_{\star};\xi)-\nabla\bar{f}(\mathbf{w}_{\star})\|^{2}$, and $G>0$ is such that $\|\nabla f(\mathbf{w};\xi)\|\leq G$ for any $\mathbf{w}$ and $\xi$. Dependence on constant numerical factors and the distance to an optimal solution is suppressed.
Setting Reference Bound at Iteration $T$ Last Iterate Guarantee Convergence for $\bm{\eta=1/\beta}$
Stochastic (*) Shamir and Zhang (2013) $\frac{1}{\eta T}+\eta G^{2}\log T$
Deterministic Smooth ($\sigma=0$) Nesterov (1998) $\frac{1}{(2-\eta\beta)\eta T}$
Stochastic Smooth Lan (2012) $\frac{1}{\eta T}+\eta\sigma^{2}$
Liu and Zhou (2024) $\frac{1}{\eta T}+\eta\sigma^{2}\log T$
Stochastic Smooth Realizable ($\sigma_{\star}=0$) Srebro et al. (2010) $\frac{1}{(1-\eta\beta)\eta T}$
Stochastic Regression Realizable ($\sigma_{\star}=0$) Bach and Moulines (2013) $\frac{1}{\eta T}$
Varre et al. (2021) $\frac{1}{(1-2\eta\beta\log T)\eta T}$
Ours (2025) $\frac{1}{(2-\eta\beta)\eta T^{1-\eta\beta(1-\eta\beta/4)}}$
(*) Shamir and Zhang (2013) consider bounded domains; Orabona (2020); Liu and Zhou (2024) obtain similar bounds for the unconstrained case. For non-fixed step sizes Jain et al. (2019) obtain minimax optimal bounds without log factors.
Our main result establishes last-iterate guarantees for with-replacement SGD, defined next. Given an initialization $\mathbf{w}_{0}\in\mathbb{R}^{d}$ and step-size $\eta>0$:
$\mathbf{w}_{t+1}\leftarrow\mathbf{w}_{t}-\eta\nabla f(\mathbf{w}_{t};i_{t}),\quad i_{t}\sim\mathcal{D}.$ (2)
Below, we state our theorem and then provide an overview of the analysis.
Theorem 10 (Last-Iterate Bound for Realizable Regression).
Consider the $\beta$-smooth, realizable Setup 1. Then, for any initialization $\mathbf{w}_{0}\in\mathbb{R}^{d}$, with-replacement SGD (Eq. 2) with step size $\eta<2/\beta$, holds:
$\mathbb{E}\bar{f}(\mathbf{w}_{T})\leq\frac{eD^{2}}{2\eta(2-\eta\beta)T^{1-\eta\beta\left(1-\eta\beta/4\right)}}\,,\quad\forall T\geq 1\,,$
where $D\triangleq\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|$. In particular, for $\eta=\frac{1}{\beta}$:
$\mathbb{E}\bar{f}(\mathbf{w}_{T})\leq\frac{e\beta D^{2}}{2\sqrt[4]{T}}.$
The important feature of Theorem 10 is the $(2-\eta\beta)$ factor in the denominator, replacing the common $(1-\eta\beta)$ of the standard analysis. This difference makes our theorem applicable to the continual regression setting which requires setting $\eta=1/\beta$ (Reduction 2). In addition, for $\eta={1}/{\left({\beta\log T}\right)}$, we recover the near-optimal rate obtained by Varre et al. (2021), i.e., $\mathbb{E}\bar{f}(\mathbf{w}_{T})=\mathcal{O}\left({\frac{\beta D^{2}\log T}{T}}\right).$
5.1 Analysis overview
In this section, we briefly outline the proof of Theorem 10, which follows immediately by combining the two lemmas below (while noting that $\eta<2/\beta\Rightarrow e^{\eta\beta\left(1-\eta\beta/4\right)}\leq e$). The first step of the proof is to establish a regret bound for SGD when applied to $f(\mathbf{w};i_{1})\ldots f(\mathbf{w};i_{T})$, holding for any step size $\eta<2/\beta$. This already departs from the standard $\eta<1/\beta$ mandated by standard analysis. All proofs for this section are given in Section E.1.
Lemma 11 (Gradient Descent Regret Bound for Smooth Optimization).
Consider the $\beta$-smooth, realizable Setup 1, and let $T\geq 1$, $(i_{0},\ldots,i_{T})\in\mathcal{I}^{T+1}$ be an arbitrary sequence of indices in $\mathcal{I}$, and ${\mathbf{w}_{0}\in\mathbb{R}^{d}}$ be an arbitrary initialization. Then, the gradient descent iterates given by ${\mathbf{w}_{t+1}\leftarrow\mathbf{w}_{t}-\eta\nabla f(\mathbf{w}_{t};i_{t})}$ for a step size $\eta<2/\beta$, hold:
$\sum_{t=0}^{T}f\left(\mathbf{w}_{t};i_{t}\right)\leq\frac{\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{2\eta(2-\eta\beta)}\,.$
With the above lemma in place, the second and main step of the analysis is to relate the loss of the last SGD iterate to the regret of the algorithm. For this, we carefully adapt an existing approach for last-iterate convergence in the non-smooth case (Shamir and Zhang, 2013). The result, given below, is slightly more general to accommodate without-replacement sampling, which we address in the next section.
Lemma 12.
Consider the $\beta$-smooth, realizable Setup 1. Let $T\geq 1$. Assume $\mathcal{P}$ is a distribution over $\mathcal{I}^{T+1}$ such that for every $0\leq t\leq\tau_{1}\leq\tau_{2}\leq T$, the following holds: For any ${i_{0},\ldots i_{t-1}\in\mathcal{I}^{t}},i\in\mathcal{I}$, $\Pr(i_{\tau_{1}}=i|i_{0},\ldots,i_{t-1})=\Pr(i_{\tau_{2}}=i|i_{0},\ldots,i_{t-1})$. Then, for any initialization $\mathbf{w}_{0}\in\mathbb{R}^{d}$, with-replacement SGD (Eq. 2) with step-size $\eta<2/\beta$, holds:
$\mathbb{E}f(\mathbf{w}_{T},i_{T})\leq(eT)^{\eta\beta\left(1-\eta\beta/4\right)}\mathbb{E}\left[\tfrac{1}{T+1}{\sum}_{t=0}^{T}f(\mathbf{w}_{t};i_{t})\right],$
where the expectation is taken with respect to $i_{0},\ldots,i_{T}$ sampled from $\mathcal{P}$.
5.2 Extension to Without-Replacement SGD
Here, we extend Theorem 10 to the without-replacement setting. Specifically, we assume =$\mathcal{I}=\{0,\ldots,n-1\}$, $\mathcal{D}=\textrm{Unif}(\mathcal{I})$, and consider gradient descent w.r.t. a uniformly random permutation of the training examples. For any initialization $\mathbf{w}_{0}\in\mathbb{R}^{d}$, step size $\eta>0$, and $\pi_{t}\sim\textrm{Unif}(\mathcal{I})$ sampled without replacement:
$\mathbf{w}_{t+1}\leftarrow\mathbf{w}_{t}-\eta\nabla f(\mathbf{w}_{t};\pi_{t})\,,$ (3)
where $f(\mathbf{w};i)\triangleq\frac{1}{2}\left\|{\mathbf{A}_{i}\mathbf{w}-\mathbf{b}_{i}}\right\|^{2}$ as defined in Setup 1. Our main result is given below.
Theorem 13 (Without-Replacement Bound).
Consider the $\beta$-smooth, realizable Setup 1. Define for all $T\!\geq\!2$, $\hat{f}_{0:T}(\mathbf{w})\!\triangleq\!\frac{1}{T+1}\sum_{t=0}^{T}f(\mathbf{w};\pi_{t})$. Then, without-replacement SGD (Eq. (3)) with step-size $\eta<{2}/{\beta}$, holds:
$\mathbb{E}_{\pi}\hat{f}_{0:T}(\mathbf{w}_{T})\leq\frac{eD^{2}}{\eta(2-\eta\beta)T^{1-\eta\beta\left(1-\eta\beta/4\right)}}+\frac{4\beta^{2}\eta D^{2}}{T}\,,\quad\forall T=2,\dots,n-1\,,$
where $D\triangleq\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|$. In particular,
$\eta=\frac{1}{\beta\log T}\Longrightarrow\mathbb{E}\hat{f}_{0:T-1}(\mathbf{w}_{T})\leq{\displaystyle\frac{14\beta D^{2}\log T}{T}}\,,\quad\quad\eta=\frac{1}{\beta}\Longrightarrow\mathbb{E}\hat{f}_{0:T-1}(\mathbf{w}_{T})\leq\frac{7\beta D^{2}}{\sqrt[4]{T}}\,.$
The proof, given in Section E.2, is based on the algorithmic stability of SGD (Bousquet and Elisseeff, 2002; Shalev-Shwartz et al., 2010; Hardt et al., 2016), and more specifically, on a variant of stability, suitable for without replacement sampling (Sherman et al., 2021; Koren et al., 2022).
6 Extensions
6.1 A Universal $\mathcal{O}(1/\sqrt[4]{k})$ Rate for General Projections Onto Convex Sets
Projections Onto Convex Sets (POCS) is a classical method that iteratively projects onto closed convex sets to find a point in their intersection (Gubin et al., 1967; Boyd et al., 2003). Formally,
Scheme 4 Projections onto Convex Sets (POCS)
  Input: A set of $T$ closed convex sets $\mathcal{C}_{1},\dots,\mathcal{C}_{T}$; an initial $\mathbf{w}_{0}\in\mathbb{R}^{d}$; an ordering $\tau:\left[{k}\right]\to\left[{T}\right]$
   For each iteration $t=1,\dots,k$:
       $\mathbf{w}_{t}\leftarrow{\Pi}_{\tau(t)}(\mathbf{w}_{t-1})\triangleq\operatorname*{argmin}_{\mathbf{w}\in\mathcal{C}_{\tau(t)}}\left\|{\mathbf{w}-\mathbf{w}_{t-1}}\right\|$
[…]
Reduction 3 (POCS $\Rightarrow$ SGD).
Consider $T$ arbitrary (nonempty) closed convex sets $\mathcal{C}_{1},\dots,\mathcal{C}_{T}$, initial point $\mathbf{w}_{0}\in\mathbb{R}^{d}$, and ordering $\tau$. Define $f_{m}(\mathbf{w})=\frac{1}{2}\left\|{\mathbf{w}-{\Pi}_{m}(\mathbf{w})}\right\|^{2},\forall m\in\left[{T}\right]$. Then,
  1. (i)
    $f_{m}$ is convex and $1$-smooth.
  2. (ii)
    The POCS update is equivalent to an SGD step: $\mathbf{w}_{t}={\Pi}_{\tau(t)}(\mathbf{w}_{t-1})=\mathbf{w}_{t-1}-\nabla_{\mathbf{w}}f_{\tau(t)}(\mathbf{w}_{t-1})$.
We can now employ our analysis from Section 5 to yield a universal rate.
Theorem 14 (Universal POCS Rate).
Consider the same conditions of Reduction 3 and assume a nonempty set intersection $\mathcal{C}_{\star}=\bigcap_{m=1}^{T}\mathcal{C}_{m}\neq\varnothing$. Then, under a random ordering with or without replacement, the expected “residual” of Scheme 4 after $\forall k\geq 1$ iterations (without replacement: $k\in\left[{T}\right]$) is bounded as,
$\mathbb{E}_{\tau}\Big{[}\frac{1}{2T}{\sum}_{m=1}^{T}\left\|{\mathbf{w}_{k}-{\Pi}_{m}(\mathbf{w}_{k})}\right\|^{2}\Big{]}=\mathbb{E}_{\tau}\Big{[}\frac{1}{2T}{\sum}_{m=1}^{T}\mathrm{dist}^{2}(\mathbf{w}_{k},\mathcal{C}_{m})\Big{]}\leq\frac{7}{\sqrt[4]{k}}\min_{\mathbf{w}\in\mathcal{C}_{\star}}\left\|{\mathbf{w}_{0}-\mathbf{w}}\right\|^{2}\,.$
To the best of our knowledge, the above is the first universal rate in the POCS literature, independent of problem parameters such as regularity or complexity measure, as demonstrated in the next Section 6.2. Universal rates are only achievable when analyzing individual distances, i.e., $f_{m}(w)=\left\|{\mathbf{w}-{\Pi}_{m}(\mathbf{w})}\right\|^{2}=\mathrm{dist}^{2}(\mathbf{w},\mathcal{C}_{m})$, rather than the distance to the intersection, i.e., $\mathrm{dist}^{2}(\mathbf{w},\mathcal{C}_{\star})$. In machine learning, the squared distance from individual sets is linked to important losses like MSE in regression or squared hinge loss in classification (Evron et al., 2022, 2023), naturally leading us to our next continual learning model.
6.2 A Universal $\mathcal{O}(1/\sqrt[4]{k})$ Rate for Random Orderings in Continual Linear Classification
Regularization methods are commonly used to prevent forgetting in continual learning (see Kirkpatrick et al., 2017). Evron et al. (2023) studied a regularized linear model for continual classification. They considered $T\geq 2$ jointly separable, binary classification tasks, defined by datasets $S_{1},\dots,S_{T}$ consisting of vectors $\mathbf{x}\in\mathbb{R}^{d}$ and their labels $y\in\left\{-1,+1\right\}$. They proved that a weakly-regularized scheme implicitly applies sequential max-margin projections. That is, in the limit as $\lambda\to 0$, the iterates of the two following schemes align in direction, enabling the study of continual classification through the lens of projection algorithms.
Scheme 5 Regularized Continual Classification
  Initialize $\mathbf{w}_{0}^{(\lambda)}=\mathbf{0}_{d}$
   For each task $t=1,\dots,k$:
       $\mathbf{w}_{t}^{(\lambda)}\leftarrow\operatorname*{argmin}_{\mathbf{w}\in\mathbb{R}^{d}}\!\!\!\!{\sum_{\,\,\,(\mathbf{x},y)\in S_{t}}\!\!\!\!e^{-y\mathbf{w}^{\top}\mathbf{x}}+\frac{\lambda}{2}\,\big{\|}{\mathbf{w}-\mathbf{w}_{t-1}^{(\lambda)}}\big{\|}^{2}}$
Scheme 6 Sequential Max-Margin Projections
  Initialize $\mathbf{w}_{0}=\mathbf{0}_{d}$
   For each task $t=1,\dots,k$:
       $\mathbf{w}_{t}\leftarrow{\Pi}_{\tau(t)}(\mathbf{w}_{t-1})\triangleq\operatorname*{argmin}_{\mathbf{w}\in\mathcal{C}_{\tau(t)}}\left\|{\mathbf{w}-\mathbf{w}_{t-1}}\right\|$
       where $\mathcal{C}_{m}\!\triangleq\!\left\{\mathbf{w}\in\mathbb{R}^{d}\mid y\mathbf{w}^{\top}\mathbf{x}\geq 1,~{}\forall(\mathbf{x},y)\!\in\!S_{m}\right\}$
They studied forgetting under several orderings, using an equivalent of our Definition 3:
$F_{\tau}(k)=\frac{1}{k}\sum_{t=1}^{k}\big{(}{\mathcal{L}_{\tau(t)}(\mathbf{w}_{k})-\mathcal{L}_{\tau(t)}(\mathbf{w}_{t})}\big{)}\leq\frac{R^{2}}{2k}\sum_{t=1}^{k}\left\|{\mathbf{w}_{k-1}-{\Pi}_{\tau(t)}(\mathbf{w}_{k-1})}\right\|^{2}\,.$
Our POCS rate (Theorem 14) combined with SGD stability arguments give the following.
Theorem 15.
Under a random ordering, with or without replacement, over $T$ jointly separable tasks, the expected forgetting of the weakly-regularized Scheme 5 (at $\lambda\to 0$) after $k\geq 1$ iterations is bounded as
$\mathbb{E}_{\tau}\big{[}F_{\tau}({k})\big{]}\leq\frac{7\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{\sqrt[4]{k}}\,,\quad\text{where }\mathbf{w}_{\star}\triangleq{\min}_{\mathbf{w}\in\mathcal{C}_{1}\cap\dots\cap\mathcal{C}_{T}}\left\|{\mathbf{w}_{0}-\mathbf{w}}\right\|^{2}\,.$
As shown in Table 3, our rate is universal while the previous one is highly dependent on $\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}$, often thought of as the “complexity” of classification problems. For example, after $k=4T\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}$ iterations, it gives a $e^{-1}$ (normalized) forgetting rate while our rate gives a potentially much smaller $\frac{5}{T^{1/4}\sqrt{\left\|\mathbf{w}_{\star}\right\|R}}$.
Table 3: Forgetting Rates in Weakly-Regularized Continual Linear Classification on Separable Data. In all cells, we omit mild constant multiplicative factors and normalize by an unavoidable $\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}$ term.
Paper / Ordering Random with Replacement Random w/o Replacement Cyclic
Evron et al. (2023) $\exp\Big{(}{-\frac{k}{4T\|{\mathbf{w}_{\star}}\|^{2}R^{2}}}\Big{)}$ $\frac{T^{2}}{\sqrt{k}}\wedge\exp\Big{(}{-\frac{k}{16T^{2}\|{\mathbf{w}_{\star}}\|^{2}R^{2}}}\Big{)}$
Ours (2025) $\frac{1}{\sqrt[4]{k}}$ $\frac{1}{\sqrt[4]{T}}$
7 Discussion
Our work established a reduction from continual linear regression to the (block) Kaczmarz method and then to “stepwise-optimal” SGD. This enabled the development of analytic tools for last-iterate SGD schemes, leading to significantly improved and even universal rates for random orderings in continual learning and the Kaczmarz method. Our main results are summarized in Tables 1 and 2.
Much of the related work has been covered throughout the paper. A further discussion of related work can be found in Appendix A. Here, we briefly highlight additional aspects of our work.
Random Continual Benchmarks.
Many popular continual benchmarks in deep learning implicitly assume a random ordering, such as the permuted MNIST benchmark (Kirkpatrick et al., 2017). In this paper, we showed that in sufficiently long task sequences, random ordering is enough to prevent catastrophic forgetting, and the training loss goes to zero, even in the worst case. In accordance with our results, Lesort et al. (2023) examined a continual learning random benchmark—in which a subset of classes is randomly sampled in each task—and observed that forgetting diminishes as the number of sampled tasks increases, even when training with standard SGD (without any modifications to mitigate forgetting). This suggests that random orderings may contaminate continual learning benchmarks, making it harder to isolate the algorithmic effects being tested. Furthermore, real-world tasks often change gradually, not adhering to random orderings. Such “gradually evolving” datasets might be more challenging and perhaps more relevant as continual benchmarks.
Connections to the Kaczmarz Method.
In Section 3.1 we revisited known connections between continual regression and the Kaczmarz method (Evron et al., 2022). We broadened this connection in Section 3.2, bridging the block Kaczmarz method and “stepwise-optimal” SGD, thus applying our novel SGD bounds to the Kaczmarz method. Using Kaczmarz terminology, given a system $\mathbf{A}\mathbf{x}=\mathbf{b}$ consisting of $T$ blocks of an average rank $\bar{r}$ where $\mathbf{A}_{m}\in\mathbb{R}^{{n_{m}}\times d},\mathbf{b}_{m}\in\mathbb{R}^{n_{m}}$, our rates from Section 4 can be summarized as $\mathbb{E}_{\tau}\big{[}\frac{1}{2T}\sum_{m=1}^{T}\big{\|}{\mathbf{A}_{m}\mathbf{x}_{k}-\mathbf{b}_{m}}\big{\|}^{2}\big{]}=\mathcal{O}\big{(}{\min\big{(}{k^{-1/4},\frac{1}{k}\sqrt{d-\bar{r}},\frac{1}{k}\sqrt{T\bar{r}}}\big{)}}\big{)}$ for random orderings with replacement and $\mathcal{O}\big{(}{\min\big{(}{k^{-1/4},\frac{1}{k}\left({d-\bar{r}}\right)}\big{)}}\big{)}$ without replacement. Note that we bounded the loss, rather than the “error” $\left\|{\mathbf{w}_{k}-\mathbf{w}_{\star}}\right\|^{2}$, thus enabling the derivation of rates independent of quantities like the condition number that can make convergence arbitrarily slow.
Non-uniform Sampling.
The seminal work of Strohmer and Vershynin (2009) proposed a Kaczmarz method where rows are sampled with probability proportional to their squared norm. Our approach accommodates non-uniform sampling, including the norm-based one, which leads to a tighter version of Theorem 8, replacing the dependence on the maximum row norm $R$ with the average one. In the block version, both uniform and non-uniform variants exist (Needell and Tropp, 2014; Gower and Richtárik, 2015). As in rank-$1$, our approach should yield better bounds when weighting samples proportional to block norms, again improving the dependence on the maximum norm to the average one.
Future Work.
We narrowed the gap between existing lower and upper worst-case bounds for random orderings in continual linear regression (see Table 1). However, a considerable gap remains between $\Omega(1/k)$ and $\mathcal{O}({1/k^{1/4}})$. Generally, we conjecture that the last-iterate SGD rates can be improved beyond those in Theorem 10, and that Theorem 13 can be extended to the multi-epoch setup. Following our reductions (Section 3), improved rates for “stepwise-optimal” SGD rates would immediately refine the bounds for continual linear regression and classification.
Acknowledgments.
We thank Edgar Dobriban (University of Pennsylvania) and Amit Attia (Tel Aviv University) for fruitful discussions.
The research of DS was funded by the European Union (ERC, A-B-C-Deep, 101039436). Views and opinions expressed are however those of the author only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency (ERCEA). Neither the European Union nor the granting authority can be held responsible for them. DS also acknowledges the support of the Schmidt Career Advancement Chair in AI.
The research of TK has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 101078075). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. This work received additional support from the Israel Science Foundation (ISF, grant number 3174/23), and a grant from the Tel Aviv University Center for AI and Data Science (TAD).
NS was supported in part by the Simons Foundation and the NSF-TRIPOD Institute on Data Economics and Learning (IDEAL).
Appendix A Related Work
Most of the related work is already discussed in the main body of the paper. Here, we elaborate on several interesting connections that remain open.
Last-iterate Guarantees for SGD.
For the general (non-realizable) smooth stochastic setup, the recent work of Liu and Zhou (2024) was the first (and only, to our knowledge) to provide upper bounds on the convergence rate of the last SGD iterate. While their bounds are applicable in the realizable setting, they require non-constant step sizes to obtain non-trivial convergence, and are therefore not useful for our purposes (see Table 2). Our analysis technique in Section 5.1 borrows from the work of Shamir and Zhang (2013, also mentioned in Table 2) which, in fact, belongs to the comparatively-richer line of work on the non-smooth setting (Shamir and Zhang, 2013; Jain et al., 2019; Zamani and Glineur, 2023; Liu and Zhou, 2024). Notably, SGD in a stochastic non-realizable (either smooth or non-smooth) setup requires uniformly bounded noise assumptions, and generally cannot accommodate a constant step size independent of the optimization horizon.
Our analysis for SGD without-replacement is related to a long line of work primarily focused on the average iterate convergence rates (e.g., Recht and Ré, 2012b; Nagaraj et al., 2019; Safran and Shamir, 2020; Rajput et al., 2020; Mishchenko et al., 2020; Cha et al., 2023; Cai et al., 2023). For the non-strongly convex case, near-optimal bounds (for the average iterate) have been established for the general smooth case (Nagaraj et al., 2019; Mishchenko et al., 2020). In a subsequent work, Cai et al. (2023) refined the dependence on problem parameters for the smooth realizable case (among others). Guarantees for the last iterate have only been established recently by Cai and Diakonikolas (2025). However, their bounds decay with the number of epochs rather than the number of iterations and apply only to non-constant step sizes, making them inapplicable to our setting. Specifically, in a realizable $\beta$-smooth setup, after $J$ without-replacement SGD epochs over a finite sum of size $n$, Mishchenko et al. (2020); Cai et al. (2023) obtained an $O(\beta/J)$ bound for the average iterate with step size $\eta=1/(\beta n)$; and Cai and Diakonikolas (2025) derived a similar bound for the last iterate up to logarithmic factors.
Another line of work related to ours studies algorithmic stability (Bousquet and Elisseeff, 2002; Shalev-Shwartz et al., 2010) of gradient methods, which is the main technique we use in the proof of Theorem 13. Our approach is similar in nature to that of Nagaraj et al. (2019); Sherman et al. (2021); Koren et al. (2022) and primarily builds on Sherman et al. (2021), who were the first to formally introduce the notion of without-replacement stability. For with-replacement SGD, Hardt et al. (2016) discussed its algorithmic stability under smooth loss functions. Later, Lei and Ying (2020), improved this bound in the realizable loss case. The case we consider—i.e., the stability of without-replacement SGD under smooth and realizable loss functions—is not covered in the existing literature.
With versus Without Replacement in Kaczmarz Methods.
Our results in Section 4 establish universal bounds for random orderings, both with and without replacement. Both the with- and without-replacement variants converge linearly towards the minimum-norm solution $\mathbf{w}_{\star}$ (Gower and Richtárik, 2015; Han and Xie, 2024), but as we explained in Section 7, the rates can be arbitrarily slow. Recht and Ré (2012a) formulated a noncommutative analog of the arithmetic-geometric mean inequality that, if true, could have shown that without-replacement orderings lead to faster loss convergence than with-replacement orderings in Kaczmarz methods, and consequently in continual linear regression. Years later, Lai and Lim (2020) proved that this inequality does not hold in general (see also De Sa, 2020). Moreover, as in other areas, empirical studies found that row shuffling followed by cyclic orderings performs as well as i.i.d. orderings (Oswald and Zhou, 2015). This naturally connects to interesting observations and open questions regarding various forms of shuffled SGD (Bottou, 2009; Yun et al., 2021). Our rates are similar for both with- and without-replacement orderings (up to small constants), meaning they do not indicate a clear advantage for either. However, we believe they are far from tight, leaving interesting open questions in this direction.
Connections to Normalized Least Mean Squares.
The NLMS algorithm is a classical adaptive filtering method. In its simplest version (Slock, 1993), the method perfectly fits a single—usually noisy—random sample at a time, using the same update rule as the Kaczmarz method (and thus, as our continual Scheme 1 in a rank-$1$ case). There also exists a more complex version of this method, which uses more samples per update (Sankaran and Beex, 2000). Both papers give strong $\mathcal{O}(1/k)$ MSE rates in the noiseless setting (matching our realizable setting). However, they assume a very limited data model, where the sampled vectors are either orthogonal or identical up-to-scaling. Under such conditions, Evron et al. (2022) showed that there is no forgetting (of previously learned tasks), implying that the MSE decays as the number of tasks still unseen at time $k$.
Appendix B Auxiliary Proofs
Lemma 16 (Bounding Forgetting Using the Training Loss).
In a realizable setting (Assumption 1), the iterates of Scheme 1 under a random task ordering $\tau$ (with or without replacement) hold $\forall k\geq 1$,
$\mathbb{E}_{\tau}[F_{\tau}(k)]$ $=\mathbb{E}_{\tau}\bigg{[}\frac{1}{2k}\sum_{t=1}^{k}\left\|{\mathbf{X}_{\tau(t)}\mathbf{w}_{k}-\mathbf{y}_{\tau(t)}}\right\|^{2}\bigg{]}\leq\mathbb{E}_{\tau}\left\|\mathbf{X}_{\tau\left(k\right)}\mathbf{w}_{k-1}-\mathbf{y}_{\tau\left(k\right)}\right\|^{2}+\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}\,,$
where $R\triangleq\max_{m\in\left[{T}\right]}\left\|{\mathbf{X}_{m}}\right\|$ is the “radius” of the data. Notice that the dependence of $\mathbf{w}_{k-1}$ on $\tau_{1},\dots,\tau_{k-1}$ is implicit. Particularly, in an ordering with replacement, we get,
$\mathbb{E}_{\tau}[F_{\tau}(k)]$ $\leq\mathbb{E}_{\tau}\bigg{[}\frac{1}{T}\sum_{m=1}^{T}\left\|{\mathbf{X}_{m}\mathbf{w}_{k-1}-\mathbf{y}_{m}}\right\|^{2}\bigg{]}+\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}=2\mathbb{E}_{\tau}\left[\mathcal{L}\left(\mathbf{w}_{k-1}\right)\right]+\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}\,.$
Proof.
As discussed in Section 3.1, Scheme 2 governs the updates of the iterates $\mathbf{w}_{t}\in\mathbb{R}^{d}$. Under Assumption 1, we define the orthogonal projection as ${\mathbf{P}}_{\tau(t)}\triangleq\mathbf{I}_{d}-\mathbf{X}_{\tau(t)}^{+}\mathbf{X}_{\tau(t)}$, revealing a recursive form:
$\mathbf{w}_{t}$ $=\mathbf{X}_{\tau(t)}^{+}\mathbf{y}_{\tau(t)}+\left(\mathbf{I}_{d}-\mathbf{X}_{\tau(t)}^{+}\mathbf{X}_{\tau(t)}\right)\mathbf{w}_{t-1}$
[Assumption 1]delimited-[]Assumption $\end{subarray}\right]$ $=\mathbf{X}_{\tau(t)}^{+}\mathbf{X}_{\tau(t)}\mathbf{w}_{\star}+\left(\mathbf{I}_{d}-\mathbf{X}_{\tau(t)}^{+}\mathbf{X}_{\tau(t)}\right)\mathbf{w}_{t-1}=(\mathbf{I}_{d}-{\mathbf{P}}_{\tau\left({t}\right)})\mathbf{w}_{\star}+{\mathbf{P}}_{\tau\left({t}\right)}\mathbf{w}_{t-1}$
$\mathbf{w}_{t}-\mathbf{w}_{\star}$ $={\mathbf{P}}_{\tau\left({t}\right)}\left({\mathbf{w}_{t-1}-\mathbf{w}_{\star}}\right)$ (4)
$\mathbf{w}_{t}-\mathbf{w}_{\star}$ $={\mathbf{P}}_{\tau\left(t\right)}\cdots{\mathbf{P}}_{\tau\left(1\right)}\left(\mathbf{w}_{0}-\mathbf{w}_{\star}\right)\,.$ (5)
We show that,
$\mathbb{E}_{\tau}\left[F_{\tau}\left(k\right)\right]$ $=\frac{1}{2k}\sum_{t=1}^{k}\mathbb{E}_{\tau}\left\|\mathbf{X}_{\tau\left(t\right)}\mathbf{w}_{k}-\mathbf{y}_{\tau\left(t\right)}\right\|^{2}=\frac{1}{2k}\sum_{t=1}^{k}\mathbb{E}_{\tau}\left\|\mathbf{X}_{\tau\left(t\right)}\left(\mathbf{w}_{k}-\mathbf{w}_{\star}\right)\right\|^{2}$
$=\frac{1}{2k}\sum_{t=1}^{k}\mathbb{E}_{\tau}\left\|\mathbf{X}_{\tau\left(t\right)}{\mathbf{P}}_{\tau\left(k\right)}\cdots{\mathbf{P}}_{\tau\left(t+1\right)}{\mathbf{P}}_{\tau\left(t\right)}\left(\mathbf{w}_{t-1}-\mathbf{w}_{\star}\right)\right\|^{2}$
$=\frac{1}{2k}\sum_{t=1}^{k}\mathbb{E}_{\tau}\left\|\mathbf{X}_{\tau\left(t\right)}{\mathbf{P}}_{\tau\left(k\right)}\cdots{\mathbf{P}}_{\tau\left(t+1\right)}\left(\mathbf{I}-{\mathbf{P}}_{\tau\left(t\right)}-\mathbf{I}\right)\left(\mathbf{w}_{t-1}-\mathbf{w}_{\star}\right)\right\|^{2}$
$\left[\begin{subarray}{c}\text{Jensen}\end{subarray}\right]$ $\leq\frac{1}{k}\sum_{t=1}^{k}\bigg{(}~{}\mathbb{E}_{\tau}\underbrace{\left\|\mathbf{X}_{\tau\left(t\right)}{\mathbf{P}}_{\tau\left(k\right)}\cdots{\mathbf{P}}_{\tau\left(t+1\right)}\left(\mathbf{I}-{\mathbf{P}}_{\tau\left(t\right)}\right)\left(\mathbf{w}_{t-1}-\mathbf{w}_{\star}\right)\right\|^{2}}_{\leq R^{2}\left\|\left(\mathbf{I}-{\mathbf{P}}_{\tau\left(t\right)}\right)\left(\mathbf{w}_{t-1}-\mathbf{w}_{\star}\right)\right\|^{2}\text{, since projections contract}}+$
$\hskip 45.00006pt\mathbb{E}_{\tau}\left\|\mathbf{X}_{\tau\left(t\right)}{\mathbf{P}}_{\tau\left(k\right)}\cdots{\mathbf{P}}_{\tau\left(t+1\right)}\left(\mathbf{w}_{t-1}-\mathbf{w}_{\star}\right)\right\|^{2}~{}\bigg{)}$
$\leq\frac{1}{k}\sum_{t=1}^{k}\bigg{(}~{}R^{2}\mathbb{E}_{\tau}\left\|\left(\mathbf{I}-{\mathbf{P}}_{\tau\left(t\right)}\right)\left(\mathbf{w}_{t-1}-\mathbf{w}_{\star}\right)\right\|^{2}+$
$\hskip 45.00006pt\mathbb{E}_{\tau}\left\|\mathbf{X}_{\tau\left(t\right)}{\mathbf{P}}_{\tau\left(k\right)}\cdots{\mathbf{P}}_{\tau\left(t+1\right)}{\mathbf{P}}_{\tau\left(t-1\right)}\cdots{\mathbf{P}}_{\tau\left(1\right)}\left(\mathbf{w}_{0}-\mathbf{w}_{\star}\right)\right\|^{2}~{}\bigg{)}\,.$
For the first term, we employ the Pythagorean theorem for orthogonal projections to get a telescoping sum and show that
$\frac{R^{2}}{k}\sum_{t=1}^{k}\mathbb{E}_{\tau}\left\|\left(\mathbf{I}-{\mathbf{P}}_{\tau\left(t\right)}\right)\left({\mathbf{w}_{t-1}-\mathbf{w}_{\star}}\right)\right\|^{2}$ $=\frac{R^{2}}{k}\sum_{t=1}^{k}\left(\mathbb{E}_{\tau}\left\|\mathbf{w}_{t-1}-\mathbf{w}_{\star}\right\|^{2}-\mathbb{E}_{\tau}\left\|{\mathbf{P}}_{\tau\left(t\right)}\left(\mathbf{w}_{t-1}-\mathbf{w}_{\star}\right)\right\|^{2}\right)$
$=\frac{R^{2}}{k}\sum_{t=1}^{k}\left(\mathbb{E}_{\tau}\left\|\mathbf{w}_{t-1}-\mathbf{w}_{\star}\right\|^{2}-\mathbb{E}_{\tau}\left\|\mathbf{w}_{t}-\mathbf{w}_{\star}\right\|^{2}\right)$
$=\frac{R^{2}}{k}\bigg{(}{\underbrace{\mathbb{E}_{\tau}\left\|\mathbf{w}_{0}-\mathbf{w}_{\star}\right\|^{2}}_{=\left\|\mathbf{w}_{\star}\right\|^{2}}-\underbrace{\mathbb{E}_{\tau}\left\|\mathbf{w}_{k}-\mathbf{w}_{\star}\right\|^{2}}_{\geq 0}}\bigg{)}\leq\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}\,.$
For the second term, we use the exchangeability of $\tau$ which applies with or without replacement,
$\mathbb{E}_{\tau}\left\|\mathbf{X}_{\tau\left(t\right)}{\mathbf{P}}_{\tau\left(k\right)}\cdots{\mathbf{P}}_{\tau\left(t+1\right)}{\mathbf{P}}_{\tau\left(t-1\right)}\cdots{\mathbf{P}}_{\tau\left(1\right)}\left(\mathbf{w}_{0}-\mathbf{w}_{\star}\right)\right\|^{2}$
$=\mathbb{E}_{\tau}\left\|\mathbf{X}_{\tau\left(k\right)}{\mathbf{P}}_{\tau\left(k-1\right)}\cdots{\mathbf{P}}_{\tau\left(1\right)}\left(\mathbf{w}_{0}-\mathbf{w}_{\star}\right)\right\|^{2}=\mathbb{E}_{\tau}\left\|\mathbf{X}_{\tau\left(k\right)}\left(\mathbf{w}_{k-1}-\mathbf{w}_{\star}\right)\right\|^{2}\,.$
Combining the two, we get
$\mathbb{E}_{\tau}\left[F_{\tau}\left(k\right)\right]$ $\leq\mathbb{E}_{\tau}\left\|\mathbf{X}_{\tau\left(k\right)}\mathbf{w}_{k-1}-\mathbf{y}_{\tau\left(k\right)}\right\|^{2}+\frac{\left\|\mathbf{w}_{\star}\right\|^{2}R^{2}}{k}\,,$
which completes the first part of the proof.
For the second part, simply notice that in an i.i.d. setting, the index $\tau(k)\sim\textrm{Unif}\left({\left[{T}\right]}\right)$ is independent of earlier indices (which yielded $\mathbf{w}_{k-1}$), and thus
$\mathbb{E}_{\tau}\left\|\mathbf{X}_{\tau\left(k\right)}\mathbf{w}_{k-1}-\mathbf{y}_{\tau\left(k\right)}\right\|^{2}=\mathbb{E}_{\tau}\bigg{[}\frac{1}{T}\sum_{m=1}^{T}\left\|{\mathbf{X}_{m}\mathbf{w}_{k-1}-\mathbf{y}_{m}}\right\|^{2}\bigg{]}\,.$
Proposition 17 (Bounding The Training Loss Using Forgetting in Without-Replacement Orderings).
Under a random ordering $\tau$ without replacement, the iterates of Scheme 1 (continual regression) satisfy $\forall k\in\left[{T}\right]$:
$\mathbb{E}_{\tau}\left[\mathcal{L}\left(\mathbf{w}_{k}\right)\right]=\frac{k}{T}\mathbb{E}_{\tau}\left[F_{\tau}\left(k\right)\right]+\frac{T-k}{2T}\mathbb{E}_{\tau}\left\|\mathbf{X}_{\tau\left(k+1\right)}\mathbf{w}_{k}-\mathbf{y}_{\tau\left(k+1\right)}\right\|^{2}\,.$
Similarly, the iterates of Scheme 4 (POCS) satisfy:
$\mathbb{E}_{\tau}\left[\mathcal{L}\left(\mathbf{w}_{k}\right)\right]=\frac{k}{T}\mathbb{E}_{\tau}\left[F_{\tau}\left(k\right)\right]+\frac{T-k}{2T}\mathbb{E}_{\tau}\left\|\mathbf{w}_{k}-{\Pi}_{\tau\left(k+1\right)}\left(\mathbf{w}_{k}\right)\right\|^{2}\,,$
where in such a POCS setting, the loss and forgetting are defined as:
$\mathcal{L}\left(\mathbf{w}_{k}\right)$ $=\frac{1}{2T}\sum_{m=1}^{T}\left\|\mathbf{w}_{k}-{\Pi}_{m}\left(\mathbf{w}_{k}\right)\right\|^{2},\quad F_{\tau}\left(k\right)=\frac{1}{2k}\sum_{t=1}^{k}\left\|\mathbf{w}_{k}-{\Pi}_{\tau\left(t\right)}\left(\mathbf{w}_{k}\right)\right\|^{2}\,.$
Proof.
We first prove the claim in the continual regression setting. If $k=T$ then $\mathbb{E}_{\tau}\left[\mathcal{L}\left(\mathbf{w}_{k}\right)\right]=\mathbb{E}_{\tau}\left[F_{\tau}\left(k\right)\right]$, and the claim follows. For $k<T$, we have:
$\mathbb{E}_{\tau}\left[\mathcal{L}\left(\mathbf{w}_{k}\right)\right]$ $=\frac{1}{2T}\sum_{m=1}^{T}\mathbb{E}_{\tau}\left\|\mathbf{X}_{m}\mathbf{w}_{k}-\mathbf{y}_{m}\right\|^{2}$
$\left[\begin{subarray}{c}\text{without replacement}\end{subarray}\right]$ $=\frac{1}{2T}\sum_{t=1}^{T}\mathbb{E}_{\tau}\left\|\mathbf{X}_{\tau\left(t\right)}\mathbf{w}_{k}-\mathbf{y}_{\tau\left(t\right)}\right\|^{2}$
$=\frac{1}{2T}\sum_{t=1}^{k}\mathbb{E}_{\tau}\left\|\mathbf{X}_{\tau\left(t\right)}\mathbf{w}_{k}-\mathbf{y}_{\tau\left(t\right)}\right\|^{2}+\frac{1}{2T}\sum_{t=k+1}^{T}\mathbb{E}_{\tau}\left\|\mathbf{X}_{\tau\left(t\right)}\mathbf{w}_{k}-\mathbf{y}_{\tau\left(t\right)}\right\|^{2}$
$=\frac{k}{T}\mathbb{E}_{\tau}\left[F_{\tau}\left(k\right)\right]+\frac{1}{2T}\sum_{t=k+1}^{T}\mathbb{E}_{\tau}\left\|\mathbf{X}_{\tau\left(t\right)}\mathbf{w}_{k}-\mathbf{y}_{\tau\left(t\right)}\right\|^{2}$
$\left[\begin{subarray}{c}\text{exchangeability}\end{subarray}\right]$ $=\frac{k}{T}\mathbb{E}_{\tau}\left[F_{\tau}\left(k\right)\right]+\frac{T-k}{2T}\mathbb{E}_{\tau}\left\|\mathbf{X}_{\tau\left(k+1\right)}\mathbf{w}_{k}-\mathbf{y}_{\tau\left(k+1\right)}\right\|^{2}\,.$
For the POCS case, simply replace $\left\|\mathbf{X}_{m}\mathbf{w}_{k}-\mathbf{y}_{m}\right\|^{2}$ with $\left\|\mathbf{w}_{k}-{\Pi}_{m}\left(\mathbf{w}_{k}\right)\right\|^{2}$.
  • Recall Lemma 6.Consider any realizable task collection such that $\mathbf{X}_{m}\mathbf{w}_{\star}=\mathbf{y}_{m},\forall m\in\left[{T}\right]$. Define $f_{m}(\mathbf{w})=\frac{1}{2}\left\|{\mathbf{X}_{m}^{+}\mathbf{X}_{m}\left({\mathbf{w}-\mathbf{w}_{\star}}\right)}\right\|^{2}$. Then, $\forall m\in\left[{T}\right],\mathbf{w}\in\mathbb{R}^{d}$
    1. (i)
      Upper bound: $\mathcal{L}_{m}(\mathbf{w})\leq R^{2}f_{m}(\mathbf{w})\triangleq\max_{m^{\prime}\in\left[{T}\right]}\left\|{\mathbf{X}_{m^{\prime}}}\right\|^{2}f_{m}$
    2. (ii)
      Gradient:       $\nabla_{\mathbf{w}}f_{m}(\mathbf{w})=\mathbf{X}_{m}^{+}\mathbf{X}_{m}\left({\mathbf{w}-\mathbf{w}_{\star}}\right)=\mathbf{X}_{m}^{+}\mathbf{X}_{m}\mathbf{w}-\mathbf{X}_{m}^{+}\mathbf{y}_{m}$
    3. (iii)
      Convexity and Smoothness: $f_{m}$ is convex and $1$-smooth.
    Proof.
    First, we use the realizability and simple norm inequalities to obtain,
    $\mathcal{L}_{m}(\mathbf{w})=\tfrac{1}{2}\left\|{\mathbf{X}_{m}\mathbf{w}-\mathbf{y}_{m}}\right\|^{2}=\tfrac{1}{2}\left\|{\mathbf{X}_{m}(\mathbf{w}-\mathbf{w}_{\star})}\right\|^{2}\leq\tfrac{\left\|{\mathbf{X}_{m}}\right\|^{2}}{2}\left\|{\mathbf{X}_{m}^{+}\mathbf{X}_{m}(\mathbf{w}-\mathbf{w}_{\star})}\right\|^{2}\leq R^{2}f(\mathbf{w})\,.$
    Since $\mathbf{X}_{m}^{+}\mathbf{X}_{m}$ is an orthogonal projection operator—and thus symmetric and idempotent—we get,
    $\nabla_{\mathbf{w}}f_{m}(\mathbf{w})=\left(\mathbf{X}_{m}^{+}\mathbf{X}_{m}\right)^{\top}{\mathbf{X}_{m}^{+}\mathbf{X}_{m}}(\mathbf{w}-\mathbf{w}_{\star})=\mathbf{X}_{m}^{+}\mathbf{X}_{m}(\mathbf{w}-\mathbf{w}_{\star})=\mathbf{X}_{m}^{+}\mathbf{X}_{m}\mathbf{w}-\mathbf{X}_{m}^{+}\mathbf{y}_{m}\,.$
    Then, the above and the fact that projection operators are non-expansive imply that $\forall\mathbf{w},\mathbf{z}\in\mathbb{R}^{d}$,
    $\left\|{\nabla_{\mathbf{w}}f_{m}(\mathbf{w})-\nabla_{\mathbf{z}}f_{m}(\mathbf{z})}\right\|=\left\|{\mathbf{X}_{m}^{+}\mathbf{X}_{m}(\mathbf{w}-\mathbf{w}_{\star}-\mathbf{z}+\mathbf{w}_{\star})}\right\|=\left\|{\mathbf{X}_{m}^{+}\mathbf{X}_{m}(\mathbf{w}-\mathbf{z})}\right\|\leq\left\|{\mathbf{w}-\mathbf{z}}\right\|\,.$
    Finally, the convexity of $f_{m}$ is immediate since $\nabla_{\mathbf{w}}^{2}f_{m}(\mathbf{w})=\mathbf{X}_{m}^{+}\mathbf{X}_{m}\succeq\mathbf{0}$.
Appendix C Proofs for Section 4.1: A Parameter-Dependent $\mathcal{O}(1/k)$ Rate
  • Recall Theorem 7.Under a random ordering with replacement over $T$ jointly realizable tasks, the expected loss and forgetting of Schemes 12 after $k\geq 3$ iterations are upper bounded as,
    $\mathbb{E}_{\tau}\left[\mathcal{L}\left(\mathbf{w}_{k}\right)\right]$ $=\mathbb{E}_{\tau}\bigg{[}\frac{1}{2T}\sum_{m=1}^{T}\left\|{\mathbf{X}_{m}\mathbf{w}_{k}-\mathbf{y}_{m}}\right\|^{2}\bigg{]}\leq\frac{\min\left({\sqrt{d-\bar{r}},\sqrt{T\bar{r}}}\right)}{2e(k-1)}\|{\mathbf{w}_{\star}}\|^{2}R^{2}$
    $\mathbb{E}_{\tau}\left[F_{\tau}(k)\right]$ $=\mathbb{E}_{\tau}\bigg{[}\frac{1}{2k}\sum_{t=1}^{k}\left\|{\mathbf{X}_{\tau(t)}\mathbf{w}_{t}-\mathbf{y}_{\tau(t)}}\right\|^{2}\bigg{]}\leq\frac{3\min\left({\sqrt{d-\bar{r}},\sqrt{T\bar{r}}}\right)}{2\left({k-2}\right)}\|{\mathbf{w}_{\star}}\|^{2}R^{2}\,,$
    where $\bar{r}\triangleq\frac{1}{T}\sum_{m\in[T]}\operatorname{rank}(\mathbf{X}_{m})$. (Recall that ${R\triangleq\max_{m\in[T]}\left\|{\mathbf{X}_{m}}\right\|}$.)
    Here, we prove the main result, followed by the necessary auxiliary corollaries and lemmas in Section C.1.
    Proof.
    We analyze the randomized block Kaczmarz algorithm for solving the linear system $\mathbf{X}\mathbf{w}=\mathbf{y}$, where the matrix and vector are partitioned into blocks as follows:
    $\mathbf{X}=\begin{pmatrix}\mathbf{X}_{1}\\ \vdots\\ \mathbf{X}_{T}\end{pmatrix},\quad\mathbf{y}=\begin{pmatrix}\mathbf{y}_{1}\\ \vdots\\ \mathbf{y}_{T}\end{pmatrix}\,.$
    By defining $\mathbf{z}_{t}=\mathbf{w}_{t}-\mathbf{w}_{\star}$ and exploiting the recursive form of Eq. 4 from the proof of Lemma 16, we obtain $\mathbf{z}_{t}={\mathbf{P}}_{\tau(t)}\mathbf{z}_{t-1}$. Note that $\mathbf{z}_{0}=\mathbf{0}_{d}-\mathbf{w}_{\star}=-\mathbf{w}_{\star}$.
    Now, define the linear map $Q:\mathbb{R}^{d\times d}\to\mathbb{R}^{d\times d}$ as
    $Q\left[{\mathbf{A}}\right]=\mathop{\mathbb{E}}_{m\sim\text{Unif}\left({\left[{T}\right]}\right)}\left[{{\mathbf{P}}_{m}\mathbf{A}{\mathbf{P}}_{m}}\right]=\frac{1}{T}\sum_{m=1}^{T}{\mathbf{P}}_{m}\mathbf{A}{\mathbf{P}}_{m}.$ (6)
    This map plays a central role in our analysis and has been studied in similar forms in prior work (Guo et al., 2022). Note that ${\mathbf{P}}_{m}$ is an orthogonal projection, which implies that it is symmetric and idempotent. Thus,
    $\mathop{\mathbb{E}}_{\tau}\!\left[{\mathbf{z}_{t+1}\mathbf{z}_{t+1}^{\top}}\right]$ $=\mathop{\mathbb{E}}_{m,\tau}\!\left[{{\mathbf{P}}_{m}\mathbf{z}_{t}\mathbf{z}_{t}^{\top}{\mathbf{P}}_{m}^{\top}}\right]=\mathop{\mathbb{E}}_{m,\tau}\!\left[{{\mathbf{P}}_{m}\mathbf{z}_{t}\mathbf{z}_{t}^{\top}{\mathbf{P}}_{m}}\right]=\mathop{\mathbb{E}}_{m}\left[{{\mathbf{P}}_{m}\mathop{\mathbb{E}}_{\tau}\left[{\mathbf{z}_{t}\mathbf{z}_{t}^{\top}}\right]{\mathbf{P}}_{m}}\right]=Q\left[{\mathop{\mathbb{E}}_{\tau}\left[{\mathbf{z}_{t}\mathbf{z}_{t}^{\top}}\right]}\right].$
    It follows that $\mathop{\mathbb{E}}_{\tau}\left[{\mathbf{z}_{t}\mathbf{z}_{t}^{\top}}\right]=Q^{t}\left[{\mathop{\mathbb{E}}_{\tau}\left[{\mathbf{z}_{0}\mathbf{z}_{0}^{\top}}\right]}\right]=Q^{t}\left[{\mathbf{z}_{0}\mathbf{z}_{0}^{\top}}\right]=Q^{t}\left[{\left({\mathbf{w}_{0}-\mathbf{w}_{\star}}\right)\left({\mathbf{w}_{0}-\mathbf{w}_{\star}}\right)^{\top}}\right]=Q^{t}\left[{\mathbf{w}_{\star}\mathbf{w}_{\star}^{\top}}\right]$, where $Q^{t}$ denotes $t$ applications of $Q$. The map $Q$ captures the evolution of the error’s second-moment under Kaczmarz updates, offering a tractable approach to analyzing the algorithm’s convergence. The expected loss at step $t$ is given by
    $\mathbb{E}_{\tau}\left[{\mathcal{L}\left({\mathbf{w}_{t}}\right)}\right]$ $=\mathop{\mathbb{E}}_{\tau}\left[{\frac{1}{2T}\sum_{i=1}^{T}\left\|{\mathbf{X}_{i}\mathbf{w}_{t}-\mathbf{y}_{i}}\right\|^{2}}\right]=\mathop{\mathbb{E}}_{\tau}\left[{\frac{1}{2T}\sum_{i=1}^{T}\left\|{\mathbf{X}_{i}\left({\mathbf{w}_{t}-\mathbf{w}_{\star}}\right)}\right\|^{2}}\right]=\mathop{\mathbb{E}}_{\tau}\left[{\frac{1}{2T}\sum_{i=1}^{T}\left\|{\mathbf{X}_{i}\mathbf{z}_{t}}\right\|^{2}}\right]$
    $=\mathop{\mathbb{E}}_{\tau}\left[{\frac{1}{2T}\left\|{\mathbf{X}\mathbf{z}_{t}}\right\|^{2}}\right]=\mathop{\mathbb{E}}_{\tau}\left[{\frac{1}{2T}\mathbf{z}_{t}^{\top}\mathbf{X}^{\top}\mathbf{X}\mathbf{z}_{t}}\right]=\mathop{\mathbb{E}}_{\tau}\left[{\operatorname{tr}\left({\frac{1}{2T}\mathbf{X}^{\top}\mathbf{X}\mathbf{z}_{t}\mathbf{z}_{t}^{\top}}\right)}\right]$
    $=\operatorname{tr}\left({\frac{1}{2T}\mathbf{X}^{\top}\mathbf{X}\mathop{\mathbb{E}}_{\tau}\left[{\mathbf{z}_{t}\mathbf{z}_{t}^{\top}}\right]}\right)=\operatorname{tr}\left({\frac{1}{2T}\mathbf{X}^{\top}\mathbf{X}Q^{t}\left[{\mathbf{w}_{\star}\mathbf{w}_{\star}^{\top}}\right]}\right).$
    We are now ready to derive the final bound. From Lemma 24, we have
    $\frac{1}{R^{2}T}\mathbf{X}^{\top}\mathbf{X}\preccurlyeq\mathbf{X}^{+}\mathbf{X}-Q\left[{\mathbf{X}^{+}\mathbf{X}}\right].$
    Additionally, by Corollary 22, $Q^{k}\left[{\mathbf{w}_{\star}\mathbf{w}_{\star}^{\top}}\right]$ is symmetric and positive semidefinite (PSD). We also note that $\frac{1}{T}\mathbf{X}^{\top}\mathbf{X}$ is symmetric PSD. The key insight from Lemma 24, combined with the trace product inequality (Lemma 23), is that it allows the expected loss to be expressed using a polynomial in $Q$. This reformulation simplifies the convergence analysis by reducing it to examining the spectral properties of $Q$. Invoking the trace product inequality, we obtain:
    $\mathbb{E}_{\tau}\left[{\mathcal{L}\left({\mathbf{w}_{k}}\right)}\right]$ $=\operatorname{tr}\left({\frac{1}{2T}\mathbf{X}^{\top}\mathbf{X}Q^{t}\left[{\mathbf{w}_{\star}\mathbf{w}_{\star}^{\top}}\right]}\right)\leq\frac{R^{2}}{2}\operatorname{tr}\left({\left({\mathbf{X}^{+}\mathbf{X}-Q\left[{\mathbf{X}^{+}\mathbf{X}}\right]}\right)Q^{k}\left[{\mathbf{w}_{\star}\mathbf{w}_{\star}^{\top}}\right]}\right)$
[…]
    $\leq\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{2}\left\|{Q^{k}\left[{\mathbf{X}^{+}\mathbf{X}-Q\left[{\mathbf{X}^{+}\mathbf{X}}\right]}\right]}\right\|_{2}=\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{2}\left\|{\big{(}{Q^{k}\left({I-Q}\right)}\big{)}\left[{\mathbf{X}^{+}\mathbf{X}}\right]}\right\|_{2}$
    $=\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{2}\left\|{\big{(}{Q^{k-1}\left({I-Q}\right)}\big{)}Q\left[{\mathbf{X}^{+}\mathbf{X}}\right]}\right\|_{2}$
    $\leq\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{2}\left\|{\big{(}{Q^{k-1}\left({I-Q}\right)}\big{)}Q\left[{\mathbf{X}^{+}\mathbf{X}}\right]}\right\|_{F}$
    $\left[\begin{subarray}{c}\text{operator norm}\end{subarray}\right]$ $\leq\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{2}\left\|{Q^{k-1}\left({I-Q}\right)}\right\|\cdot\left\|{Q\left[{\mathbf{X}^{+}\mathbf{X}}\right]}\right\|_{F}$
[…]
    To clarify, the operator norm of a linear map $H$ is defined as $\left\|{H}\right\|=\sup_{\mathbf{A}\in\mathbb{R}^{d\times d},\left\|{\mathbf{A}}\right\|_{F}=1}\left\|{H\left[{\mathbf{A}}\right]}\right\|_{F}$. The reason for switching from the spectral norm to the Frobenius norm is to enable the use of the spectral mapping theorem to bound the operator norm of $Q^{k-1}\left({I-Q}\right)$, applicable only for inner-product-based norms. We complete the proof by bounding the forgetting using the training loss (Lemma 16). That is,
    $\mathbb{E}_{\tau}[F_{\tau}(k)]$ $=\mathbb{E}_{\tau}\bigg{[}\frac{1}{2k}\sum_{t=1}^{k}\left\|{\mathbf{X}_{\tau(t)}\mathbf{w}_{t}-\mathbf{y}_{\tau(t)}}\right\|^{2}\bigg{]}\leq 2\mathbb{E}_{\tau}\left[{\mathcal{L}\left({\mathbf{w}_{k-1}}\right)}\right]+\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}$
    $\leq\frac{3\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{2\left({k-2}\right)}\min\left({\sqrt{T\bar{r}},\sqrt{d-\bar{r}}}\right)\,.$
C.1 Key Properties and Auxiliary Lemmas
Definition 18 (Positive Map).
A positive map $H:\mathbb{R}^{d\times d}\to\mathbb{R}^{d\times d}$ is a linear map that maps PSD matrices to PSD matrices. Formally, if $\mathbf{0}\preccurlyeq\mathbf{A}\in\mathbb{R}^{d\times d}$, then $\mathbf{0}\preccurlyeq H\left[{\mathbf{A}}\right]$.
Definition 19 (Symmetric Map).
A symmetric map $H:\mathbb{R}^{d\times d}\to\mathbb{R}^{d\times d}$ is a linear map that maps symmetric matrices to symmetric matrices. Formally, if $\mathbf{A}=\mathbf{A}^{\top}\in\mathbb{R}^{d\times d}$, then $H\left[{\mathbf{A}}\right]=H\left[{\mathbf{A}}\right]^{\top}$.
Corollary 20.
$Q$, defined in Eq. 6, is a positive map.
Proof.
Let $\mathbf{0}\preccurlyeq\mathbf{A}\in\mathbb{R}^{d\times d}$. Then, for all $i\in\left[{T}\right]$, $\mathbf{0}\preccurlyeq{\mathbf{P}}_{i}\mathbf{A}{\mathbf{P}}_{i}$. Meaning $Q\left[{\mathbf{A}}\right]$ is PSD as a convex combination of PSD matrices.
Corollary 21.
$Q$ is a symmetric map. Moreover, for all $\mathbf{A}\in\mathbb{R}^{d\times d}$, it satisfies $Q\left[{\mathbf{A}}\right]^{\top}=Q\left[{\mathbf{A}^{\top}}\right]$.
Proof.
Let $\mathbf{A}\in\mathbb{R}^{d\times d}$. Then,
$Q\left[{\mathbf{A}}\right]^{\top}=\frac{1}{T}\sum_{i=1}^{T}\left({{\mathbf{P}}_{i}\mathbf{A}{\mathbf{P}}_{i}}\right)^{\top}=\frac{1}{T}\sum_{i=1}^{T}{\mathbf{P}}_{i}^{\top}\mathbf{A}^{\top}{\mathbf{P}}_{i}^{\top}=\frac{1}{T}\sum_{i=1}^{T}{\mathbf{P}}_{i}\mathbf{A}^{\top}{\mathbf{P}}_{i}=Q\left[{\mathbf{A}^{\top}}\right].$
Corollary 22.
For $n\in\mathbb{N}^{+}$, the iterated application of the map $Q$, denoted $Q^{n}$, is a positive symmetric map.
Proof.
For $n=1$, given by Corollaries 20 and 21. For $n>1$, this follows trivially by induction.
Lemma 23 (Trace Product Inequality).
Let $\mathbf{A},\mathbf{B},\mathbf{C}\in\mathbb{R}^{d\times d}$ be symmetric PSD matrices such that $\mathbf{A}\preccurlyeq\mathbf{B}$. Then, $\operatorname{tr}\left({\mathbf{A}\mathbf{C}}\right)\leq\operatorname{tr}\left({\mathbf{B}\mathbf{C}}\right)$.
Proof.
Since $\mathbf{0}\preccurlyeq\mathbf{C}=\mathbf{C}^{\top}$, it has a square symmetric PSD root $\mathbf{C}^{1/2}$. Given that $\mathbf{A},\mathbf{B}$ are symmetric and $\mathbf{A}\preccurlyeq\mathbf{B}$, it follows that $\mathbf{C}^{1/2}\mathbf{A}\mathbf{C}^{1/2}\preccurlyeq\mathbf{C}^{1/2}\mathbf{B}\mathbf{C}^{1/2}$ (from Horn and Johnson, 2012, Theorem 7.7.2.a). Applying the cyclic property of the trace and using the fact that for symmetric matrices ordered in the Löwner sense, their traces are also ordered (Horn and Johnson, 2012, Corollary 7.7.4.d), we obtain
$\operatorname{tr}\left({\mathbf{A}\mathbf{C}}\right)=\operatorname{tr}\left({\mathbf{A}\mathbf{C}^{1/2}\mathbf{C}^{1/2}}\right)=\operatorname{tr}\left({\mathbf{C}^{1/2}\mathbf{A}\mathbf{C}^{1/2}}\right)\leq\operatorname{tr}\left({\mathbf{C}^{1/2}\mathbf{B}\mathbf{C}^{1/2}}\right)=\operatorname{tr}\left({\mathbf{B}\mathbf{C}}\right).$
Lemma 24.
Let $R=\max_{i\in\left[{T}\right]}{\left\|{\mathbf{X}_{i}}\right\|}$. Then, $\frac{1}{R^{2}T}\mathbf{X}^{\top}\mathbf{X}\preccurlyeq\mathbf{X}^{+}\mathbf{X}-Q\left[{\mathbf{X}^{+}\mathbf{X}}\right]$
Proof.
We perform SVD on each $\mathbf{X}_{i}=\mathbf{U}_{i}{\mathbf{\Sigma}}_{i}\mathbf{V}_{i}^{\top}$. Then,
$\frac{1}{R^{2}T}\mathbf{X}^{\top}\mathbf{X}$ $=\frac{1}{R^{2}T}\sum_{i=1}^{T}\mathbf{X}_{i}^{\top}\mathbf{X}_{i}=\frac{1}{R^{2}T}\sum_{i=1}^{T}\mathbf{V}_{i}{\mathbf{\Sigma}}_{i}^{2}\mathbf{V}_{i}^{\top}$
On the other hand:
$\mathbf{X}^{+}\mathbf{X}-Q\left[{\mathbf{X}^{+}\mathbf{X}}\right]$ $=\mathbf{X}^{+}\mathbf{X}-\frac{1}{T}\sum_{i=1}^{T}\left({\mathbf{I}-\mathbf{X}_{i}^{+}\mathbf{X}_{i}}\right)\mathbf{X}^{+}\mathbf{X}\left({\mathbf{I}-\mathbf{X}_{i}^{+}\mathbf{X}_{i}}\right)$
$=\mathbf{X}^{+}\mathbf{X}-\frac{1}{T}\sum_{i=1}^{T}\mathbf{X}^{+}\mathbf{X}-\mathbf{X}_{i}^{+}\mathbf{X}_{i}\mathbf{X}^{+}\mathbf{X}-\mathbf{X}^{+}\mathbf{X}\mathbf{X}_{i}^{+}\mathbf{X}_{i}+\mathbf{X}_{i}^{+}\mathbf{X}_{i}\mathbf{X}^{+}\mathbf{X}\mathbf{X}_{i}^{+}\mathbf{X}_{i}$
$\left[\begin{subarray}{c}\text{Im}\left({\mathbf{X}_{i}^{+}\mathbf{X}_{i}}\right)\subseteq\text{Im}\left({\mathbf{X}^{+}\mathbf{X}}\right)\end{subarray}\right]$ $=-\frac{1}{T}\sum_{i=1}^{T}-\mathbf{X}_{i}^{+}\mathbf{X}_{i}-\mathbf{X}_{i}^{+}\mathbf{X}_{i}+\mathbf{X}_{i}^{+}\mathbf{X}_{i}=\frac{1}{T}\sum_{i=1}^{T}\mathbf{X}_{i}^{+}\mathbf{X}_{i}=\frac{1}{T}\sum_{i=1}^{T}\mathbf{V}_{i}{\mathbf{\Sigma}}_{i}^{+}{\mathbf{\Sigma}}_{i}\mathbf{V}_{i}^{\top}.$
Now consider the difference:
$\left({\mathbf{X}^{+}\mathbf{X}-Q\left[{\mathbf{X}^{+}\mathbf{X}}\right]}\right)-\frac{1}{R^{2}T}\mathbf{X}^{\top}\mathbf{X}$ $=\frac{1}{T}\sum_{i=1}^{T}\mathbf{V}_{i}\left({{\mathbf{\Sigma}}_{i}^{+}{\mathbf{\Sigma}}_{i}-\frac{1}{R^{2}}{\mathbf{\Sigma}}_{i}^{2}}\right)\mathbf{V}_{i}^{\top}.$
We know that $\frac{1}{R}\left({{\mathbf{\Sigma}}_{i}}\right)_{j,j}\in[0,1]$. We analyze two cases for each diagonal entry:
  • If $\left({{\mathbf{\Sigma}}_{i}}\right)_{j,j}=0$, then $\left({{\mathbf{\Sigma}}_{i}^{+}{\mathbf{\Sigma}}_{i}-\frac{1}{R^{2}}{\mathbf{\Sigma}}_{i}^{2}}\right)_{j,j}=0$.
  • Otherwise, $\left({{\mathbf{\Sigma}}_{i}^{+}{\mathbf{\Sigma}}_{i}}\right)_{j,j}=1$, and $\frac{1}{R^{2}}\left({{\mathbf{\Sigma}}_{i}^{2}}\right)_{j,j}\leq 1$, which gives $\left({{\mathbf{\Sigma}}_{i}^{+}{\mathbf{\Sigma}}_{i}-\frac{1}{R^{2}}{\mathbf{\Sigma}}_{i}^{2}}\right)_{j,j}\geq 0$.
Thus,
$\mathbf{0}\preccurlyeq\mathbf{V}_{i}\left({{\mathbf{\Sigma}}_{i}^{+}{\mathbf{\Sigma}}_{i}-\frac{1}{R^{2}}{\mathbf{\Sigma}}_{i}^{2}}\right)\mathbf{V}_{i}^{\top}\,.$
Averaging over all $i$, we get:
$\mathbf{0}=\frac{1}{T}\sum_{i=1}^{T}\mathbf{0}$ $\preccurlyeq\frac{1}{T}\sum_{i=1}^{T}\mathbf{V}_{i}\left({{\mathbf{\Sigma}}_{i}^{+}{\mathbf{\Sigma}}_{i}-\frac{1}{R^{2}}{\mathbf{\Sigma}}_{i}^{2}}\right)\mathbf{V}_{i}^{\top}=\left({\mathbf{X}^{+}\mathbf{X}-Q\left[{\mathbf{X}^{+}\mathbf{X}}\right]}\right)-\frac{1}{R^{2}T}\mathbf{X}^{\top}\mathbf{X}$
$\frac{1}{R^{2}T}\mathbf{X}^{\top}\mathbf{X}$ $\preccurlyeq\mathbf{X}^{+}\mathbf{X}-Q\left[{\mathbf{X}^{+}\mathbf{X}}\right]\,.$
Lemma 25.
Let $\mathbf{A},\mathbf{B}\in\mathbb{R}^{d\times d}$ and $n\in\mathbb{N}^{+}$. Then, $\operatorname{tr}\left({\mathbf{A}Q^{n}\left[{\mathbf{B}}\right]}\right)=\operatorname{tr}\left({Q^{n}\left[{\mathbf{A}}\right]\mathbf{B}}\right)$.
Proof.
From the definition of $Q$ (Eq. 6),
$\operatorname{tr}\left({\mathbf{A}Q^{n}\left[{\mathbf{B}}\right]}\right)$ $=\operatorname{tr}\left({\mathbf{A}\frac{1}{T^{n}}\sum_{j_{1},\ldots,j_{n}=1}^{T}{\mathbf{P}}_{j_{1}}\cdots{\mathbf{P}}_{j_{n}}\mathbf{B}{\mathbf{P}}_{j_{n}}\cdots{\mathbf{P}}_{j_{1}}}\right)$
$\left[\begin{subarray}{c}\text{linearity}\end{subarray}\right]$ $=\frac{1}{T^{n}}\sum_{j_{1},\ldots,j_{n}=1}^{T}\operatorname{tr}\left({\mathbf{A}{\mathbf{P}}_{j_{1}}\cdots{\mathbf{P}}_{j_{n}}\mathbf{B}{\mathbf{P}}_{j_{n}}\cdots{\mathbf{P}}_{j_{1}}}\right)$
$\left[\begin{subarray}{c}\text{cyclic property}\end{subarray}\right]$ $=\frac{1}{T^{n}}\sum_{j_{1},\ldots,j_{n}=1}^{T}\operatorname{tr}\left({{\mathbf{P}}_{j_{n}}\cdots{\mathbf{P}}_{j_{1}}\mathbf{A}{\mathbf{P}}_{j_{1}}\cdots{\mathbf{P}}_{j_{n}}\mathbf{B}}\right)$
$\left[\begin{subarray}{c}\text{linearity}\end{subarray}\right]$ $=\operatorname{tr}\left({\left({\frac{1}{T^{n}}\sum_{j_{1}\ldots,j_{n}=1}^{T}{\mathbf{P}}_{j_{n}}\cdots{\mathbf{P}}_{j_{1}}\mathbf{A}{\mathbf{P}}_{j_{1}}\cdots{\mathbf{P}}_{j_{n}}}\right)\mathbf{B}}\right)$
$=\operatorname{tr}\left({Q^{n}\left[{\mathbf{A}}\right]\mathbf{B}}\right)\,.$
Proposition 26.
$Q$ is self adjoint.
Proof.
Let $\mathbf{A},\mathbf{B}\in\mathbb{R}^{d\times d}$. Then,
$\langle Q\left[{\mathbf{A}}\right],\mathbf{B}\rangle$ $=\operatorname{tr}\left({Q\left[{\mathbf{A}}\right]^{\top}\mathbf{B}}\right)=\operatorname{tr}\left({\mathbf{B}^{\top}Q\left[{\mathbf{A}}\right]}\right)$
[…]
[Corollary 21]delimited-[]Corollary $\end{subarray}\right]$ $=\operatorname{tr}\left({Q\left[{\mathbf{B}}\right]^{\top}\mathbf{A}}\right)=\operatorname{tr}\left({\mathbf{A}^{\top}Q\left[{\mathbf{B}}\right]}\right)=\langle\mathbf{A},Q\left[{\mathbf{B}}\right]\rangle\,.$
Proposition 27.
The spectrum of $Q$ is contained in the interval $[0,1]$.
Proof.
Let $\mathbf{A}\in\mathbb{R}^{d\times d}$. Then, by definition,
$\langle Q\left[{\mathbf{A}}\right],\mathbf{A}\rangle$ $=\operatorname{tr}\left({Q\left[{\mathbf{A}}\right]^{\top}\mathbf{A}}\right)=\frac{1}{T}\sum_{i=1}^{T}\operatorname{tr}\left({{\mathbf{P}}_{i}\mathbf{A}^{\top}{\mathbf{P}}_{i}\mathbf{A}}\right)$
$\left[\begin{subarray}{c}\text{idempotence,}\\ \text{cyclic property}\end{subarray}\right]$ $=\frac{1}{T}\sum_{i=1}^{T}\operatorname{tr}\left({{\mathbf{P}}_{i}\mathbf{A}^{\top}{\mathbf{P}}_{i}{\mathbf{P}}_{i}\mathbf{A}{\mathbf{P}}_{i}}\right)=\frac{1}{T}\sum_{i=1}^{T}\left\|{{\mathbf{P}}_{i}\mathbf{A}{\mathbf{P}}_{i}}\right\|_{F}^{2}\geq 0\,.$
Since each ${\mathbf{P}}_{i}$ is an orthogonal projection, its spectral norm satisfies $\left\|{{\mathbf{P}}_{i}}\right\|_{2}=1$. Applying the operator inequality $\left\|{\mathbf{X}\mathbf{Y}}\right\|_{F}\leq\left\|{\mathbf{X}}\right\|_{2}\left\|{\mathbf{Y}}\right\|_{F}$ twice, we obtain
$\frac{1}{T}\sum_{i=1}^{T}\left\|{{\mathbf{P}}_{i}\mathbf{A}{\mathbf{P}}_{i}}\right\|_{F}^{2}\leq\left\|{{\mathbf{P}}_{i}}\right\|_{2}^{4}\left\|{\mathbf{A}}\right\|_{F}^{2}=\left\|{\mathbf{A}}\right\|_{F}^{2}\,.$
Thus, for any $\mathbf{A}\in\mathbb{R}^{d\times d}$,
$0\leq\langle Q\left[{\mathbf{A}}\right],\mathbf{A}\rangle\leq\left\|{\mathbf{A}}\right\|_{F}^{2}\,.$
From the Rayleigh quotient characterization of eigenvalues, this implies that every eigenvalue $\lambda$ of $Q$ satisfies $0\leq\lambda\leq 1$, i.e., $\sigma(Q)\subset[0,1]\,$.
Lemma 28.
$\left\|{Q^{n}\left({I-Q}\right)}\right\|\leq\frac{1}{en},\quad\text{for }n\in\mathbb{N}^{+}.$
Proof.
By Proposition 26, $Q$ is self adjoint. Thus, we can apply the spectral mapping theorem to the polynomial $x\mapsto x^{n}\left({1-x}\right)$. The eigenvalues of $Q^{n}\left({I-Q}\right)$ are of the form $\lambda^{n}\left({1-\lambda}\right)$, where $\lambda$ is an eigenvalue of $Q$. From Proposition 27, we know that $\lambda\in\left[0,1\right]$. Using an algebraic property of $\lambda^{n}\left({1-\lambda}\right)$ for $\lambda\in\left[{0,1}\right]$, we conclude that $\lambda^{n}\left({1-\lambda}\right)\in\left[0,\frac{1}{en}\right]$.
Therefore, $\left\|{Q^{n}\left({I-Q}\right)}\right\|\leq\frac{1}{en}$.
Lemma 29.
$\left\|{Q\left[{\mathbf{X}^{+}\mathbf{X}}\right]}\right\|_{F}\leq\min\left({\sqrt{T\bar{r}},\sqrt{d-\bar{r}}}\right)$
Proof.
We first bound $\left\|{Q\left[{\mathbf{X}^{+}\mathbf{X}}\right]}\right\|_{F}$ using the operator norm bound on $Q$ (Proposition 27):
$\left\|{Q\left[{\mathbf{X}^{+}\mathbf{X}}\right]}\right\|_{F}$ $\leq\underbrace{\left\|{Q}\right\|}_{\leq 1}\cdot\left\|{\mathbf{X}^{+}\mathbf{X}}\right\|_{F}\leq\left\|{\mathbf{X}^{+}\mathbf{X}}\right\|_{F}=\sqrt{\operatorname{rank}\left({\mathbf{X}^{+}\mathbf{X}}\right)}=\sqrt{T\bar{r}}\,.$
Next, we use a pseudo-inverse property—that $\mathbf{X}^{+}\mathbf{X}\preccurlyeq\mathbf{I}$—and the positivity of $Q$ to show,
$\mathbf{0}$ $\preccurlyeq Q\left[{\mathbf{I}-\mathbf{X}^{+}\mathbf{X}}\right]$
$Q\left[{\mathbf{X}^{+}\mathbf{X}}\right]$ $\preccurlyeq Q\left[{\mathbf{I}}\right]$
$\left\|{Q\left[{\mathbf{X}^{+}\mathbf{X}}\right]}\right\|_{F}$ $\leq\left\|{Q\left[{\mathbf{I}}\right]}\right\|_{F}=\Big{\|}{\frac{1}{T}\sum_{i=1}^{T}{\mathbf{P}}_{i}}\Big{\|}_{F}\leq\frac{1}{T}\sum_{i=1}^{T}\left\|{{\mathbf{P}}_{i}}\right\|_{F}$
$=\frac{1}{T}\sum_{i=1}^{T}\sqrt{\operatorname{rank}\left({{\mathbf{P}}_{i}}\right)}=\frac{1}{T}\sum_{i=1}^{T}\sqrt{d-\operatorname{rank}\left({\mathbf{X}_{i}}\right)}$
$\left[\begin{subarray}{c}\text{Jensen (concave)}\end{subarray}\right]$ $\leq\sqrt{d-\bar{r}}\,.$
Appendix D Proofs of Universal Continual Regression Rates (Sections 4.2 and 4.3)
The proofs in this appendix focus on the properties of forgetting and loss, “translating” them into the language of last-iterate SGD. We then apply our last-iterate results, proved in Appendix E.
D.1 Proof of Theorem 8: A Universal $\mathcal{O}(1/\sqrt[4]{k})$ Rate
  • Recall Theorem 8.Under a random ordering with replacement over $T$ jointly realizable tasks, the expected loss and forgetting of Schemes 12 after $k\geq 2$ iterations are bounded as,
    $\mathbb{E}_{\tau}\left[\mathcal{L}\left(\mathbf{w}_{k}\right)\right]$ $=\mathbb{E}_{\tau}\bigg{[}\frac{1}{2T}\sum_{m=1}^{T}\big{\|}{\mathbf{X}_{m}\mathbf{w}_{k}-\mathbf{y}_{m}}\big{\|}^{2}\bigg{]}\leq\frac{2}{\sqrt[4]{k}}\big{\|}{\mathbf{w}_{\star}}\big{\|}^{2}R^{2}\,,$
    $\mathbb{E}_{\tau}\left[F_{\tau}(k)\right]$ $=\mathbb{E}_{\tau}\bigg{[}\frac{1}{2k}\sum_{t=1}^{k}\left\|{\mathbf{X}_{\tau(t)}\mathbf{w}_{k}-\mathbf{y}_{\tau(t)}}\right\|^{2}\bigg{]}\leq\frac{5}{\sqrt[4]{k-1}}\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}\,.$
    Proof.
    Let $\tau$ be a random with-replacement ordering, and $\mathbf{w}_{0},\ldots,\mathbf{w}_{k}$ be the corresponding iterates produced by the continual Scheme 1 (or the equivalent Kaczmarz Scheme 2). By Reduction 2, these are exactly the (stochastic) gradient descent iterates produced given an initialization $\mathbf{w}_{0}$ and a step size of $\eta=1$, on the loss sequence $f_{\tau(1)},\ldots,f_{\tau(k)}$, where we defined:
    $f_{m}(\mathbf{w})\triangleq\frac{1}{2}\left\|{\mathbf{X}_{m}^{+}\mathbf{X}_{m}(\mathbf{w}-\mathbf{w}_{\star})}\right\|^{2}.$
    Furthermore, Lemma 6 states that for all $\mathbf{w}\in\mathbb{R}^{d}$,
    $\mathcal{L}(\mathbf{w})=\frac{1}{2T}\sum_{m=1}^{T}\left\|{\mathbf{X}_{m}\mathbf{w}-\mathbf{y}_{m}}\right\|^{2}=\mathbb{E}_{m\sim\textrm{Unif}([T])}\mathcal{L}_{m}(\mathbf{w})\leq R^{2}\mathbb{E}_{m\sim\textrm{Unif}([T])}f_{m}(\mathbf{w})\,.$
    Therefore, establishing last iterate convergence of with-replacement SGD (Eq. 2) on the objective function
    $\bar{f}(\mathbf{w})\triangleq\mathbb{E}_{m\sim[T]}f_{m}(\mathbf{w})\,,$
    will imply the desired result. Indeed, again by Lemma 6, $f_{m}(\cdot)$ is $1$-smooth for all $m\in[T]$. Hence, plugging in $\mathbf{A}=\mathbf{X}_{m}^{+}\mathbf{X}_{m}\Rightarrow\left\|{\mathbf{A}}\right\|=1=\beta$ into Theorem 10, SGD with $\eta=1$ guarantees that after $k\geq 1$ gradient steps:
[…]
    and therefore $\mathbb{E}\mathcal{L}(\mathbf{w}_{k})\leq\frac{2R^{2}\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{\sqrt[4]{k}}$, which proves the first claim. The second claim follows immediately from Lemma 16, and we are done.
D.2 Proving Theorem 9: Main Result for Without Replacement Orderings
  • Recall Theorem 9.Under a random ordering without replacement over $T$ jointly realizable tasks, the expected loss and forgetting of Schemes 12 after $k\in\{2,\dots,T\}$ iterations are both bounded as,
    $\mathbb{E}\left[\mathcal{L}\left({\mathbf{w}_{k}}\right)\right],\,\mathbb{E}\left[F_{\tau}(k)\right]\leq\min\left(\frac{7}{\sqrt[4]{k-1}},~{}\frac{d-\bar{r}+1}{k-1}\right)\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}\,.$
    Proof.
From Lemmas 16 and 6, we have
$\mathop{\mathbb{E}}_{\tau}\left[{F_{\tau}\left({k}\right)}\right]\leq\mathop{\mathbb{E}}_{\tau}\left\|{\mathbf{X}_{\tau\left({k}\right)}\mathbf{w}_{k-1}-\mathbf{y}_{\tau\left({k}\right)}}\right\|^{2}+\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}\leq 2R^{2}\mathop{\mathbb{E}}_{\tau}f_{\tau(k)}(\mathbf{w}_{k-1})+\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}\,.$
Combining with Proposition 17, we get,
$\mathbb{E}_{\tau}\left[\mathcal{L}\left(\mathbf{w}_{k}\right)\right]$ $=\frac{k}{T}\mathbb{E}_{\tau}\left[F_{\tau}\left(k\right)\right]+\frac{T-k}{2T}\mathbb{E}_{\tau}\left\|\mathbf{X}_{\tau\left(k+1\right)}\mathbf{w}_{k}-\mathbf{y}_{\tau\left(t\right)}\right\|^{2}$
$\leq\frac{k}{T}\left({2R^{2}\mathop{\mathbb{E}}_{\tau}f_{\tau(k)}(\mathbf{w}_{k-1})+\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}}\right)+\frac{T-k}{2T}\mathop{\mathbb{E}}_{\tau}f_{\tau(k+1)}(\mathbf{w}_{k})$
$\left[\begin{subarray}{c}k\leq T\end{subarray}\right]$ $\leq R^{2}\left({\frac{2k}{T}\mathop{\mathbb{E}}_{\tau}f_{\tau(k)}(\mathbf{w}_{k-1})+\frac{T-k}{T}\mathop{\mathbb{E}}_{\tau}f_{\tau(k+1)}(\mathbf{w}_{k})}\right)+\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}\,.$
Thus, to bound both the expected forgetting and loss, we need to bound expressions like $\mathop{\mathbb{E}}_{\tau}\!f_{\tau(k+1)}(\mathbf{w}_{k})$.
We first prove the dimension dependent term. Note that,
$2\mathop{\mathbb{E}}_{\tau}f_{\tau(k)}(\mathbf{w}_{k-1})=\mathop{\mathbb{E}}_{\tau}\left\|{\mathbf{X}_{\tau\left({k}\right)}^{+}\mathbf{X}_{\tau\left({k}\right)}\left({\mathbf{w}_{k-1}-\mathbf{w}_{\star}}\right)}\right\|^{2}\triangleq\mathop{\mathbb{E}}_{\tau}\left\|{\left({\mathbf{I}-{\mathbf{P}}_{\tau\left({k}\right)}}\right)\left({\mathbf{w}_{k-1}-\mathbf{w}_{\star}}\right)}\right\|^{2}.$
Recall that from Eq. 5 in the proof of Lemma 16, we have
$\left({\mathbf{w}_{k-1}-\mathbf{w}_{\star}}\right)={\mathbf{P}}_{\tau\left({k-1}\right)}\cdots{\mathbf{P}}_{\tau\left({1}\right)}\left({\mathbf{w}_{0}-\mathbf{w}_{\star}}\right)=-{\mathbf{P}}_{\tau\left({k-1}\right)}\cdots{\mathbf{P}}_{\tau\left({1}\right)}\mathbf{w}_{\star}.$
Thus, we obtain
$\mathop{\mathbb{E}}_{\tau}\left\|{\left({\mathbf{I}-{\mathbf{P}}_{\tau\left({k}\right)}}\right)\left({\mathbf{w}_{k-1}-\mathbf{w}_{\star}}\right)}\right\|^{2}=\mathop{\mathbb{E}}_{\tau}\left\|{\left({\mathbf{I}-{\mathbf{P}}_{\tau\left({k}\right)}}\right){\mathbf{P}}_{\tau\left({k-1}\right)}\cdots{\mathbf{P}}_{\tau\left({1}\right)}\mathbf{w}_{\star}}\right\|^{2}$
$\leq\mathop{\mathbb{E}}_{\tau}\left\|{\left({\mathbf{I}-{\mathbf{P}}_{\tau\left({k}\right)}}\right){\mathbf{P}}_{\tau\left({k-1}\right)}\cdots{\mathbf{P}}_{\tau\left({1}\right)}}\right\|_{2}^{2}\cdot\left\|{\mathbf{w}_{\star}}\right\|^{2}\leq\left\|{\mathbf{w}_{\star}}\right\|^{2}\mathop{\mathbb{E}}_{\tau}\left\|{\left({\mathbf{I}-{\mathbf{P}}_{\tau\left({k}\right)}}\right){\mathbf{P}}_{\tau\left({k-1}\right)}\cdots{\mathbf{P}}_{\tau\left({1}\right)}}\right\|_{F}^{2}$
$=\left\|{\mathbf{w}_{\star}}\right\|^{2}\mathop{\mathbb{E}}_{\tau}\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({1}\right)}\cdots{\mathbf{P}}_{\tau\left({k-1}\right)}\left({\mathbf{I}-{\mathbf{P}}_{\tau\left({k}\right)}}\right){\mathbf{P}}_{\tau\left({k-1}\right)}\cdots{\mathbf{P}}_{\tau\left({1}\right)}}\right).$
By exchangeability,
$\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({1}\right)}\cdots{\mathbf{P}}_{\tau\left({t-1}\right)}\left({\mathbf{I}-{\mathbf{P}}_{\tau\left({t}\right)}}\right){\mathbf{P}}_{\tau\left({t-1}\right)}\cdots{\mathbf{P}}_{\tau\left({1}\right)}}\right)=\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({t}\right)}\cdots{\mathbf{P}}_{\tau\left({2}\right)}\left({\mathbf{I}-{\mathbf{P}}_{\tau\left({1}\right)}}\right){\mathbf{P}}_{\tau\left({2}\right)}\cdots{\mathbf{P}}_{\tau\left({t}\right)}}\right).$
Let us define $a_{t}=\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({t}\right)}\cdots{\mathbf{P}}_{\tau\left({2}\right)}\left({\mathbf{I}-{\mathbf{P}}_{\tau\left({1}\right)}}\right){\mathbf{P}}_{\tau\left({2}\right)}\cdots{\mathbf{P}}_{\tau\left({t}\right)}}\right).$ Then, we have
$a_{t+1}$ $=\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({t+1}\right)}\cdots{\mathbf{P}}_{\tau\left({2}\right)}\left({\mathbf{I}-{\mathbf{P}}_{\tau\left({1}\right)}}\right){\mathbf{P}}_{\tau\left({2}\right)}\cdots{\mathbf{P}}_{\tau\left({t+1}\right)}}\right)$
$\left[\begin{subarray}{c}\text{cyclic property of trace}\end{subarray}\right]$ $=\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({t+1}\right)}^{2}{\mathbf{P}}_{\tau\left({t}\right)}\cdots{\mathbf{P}}_{\tau\left({2}\right)}\left({\mathbf{I}-{\mathbf{P}}_{\tau\left({1}\right)}}\right){\mathbf{P}}_{\tau\left({2}\right)}\cdots{\mathbf{P}}_{\tau\left({t}\right)}}\right)$
$\left[\begin{subarray}{c}\text{Von Neumann's trace inequality}\end{subarray}\right]$ $\leq\underbrace{\big{\|}{{\mathbf{P}}_{\tau\left({t+1}\right)}^{2}}\big{\|}_{2}}_{=1}\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({t}\right)}\cdots{\mathbf{P}}_{\tau\left({2}\right)}\left({\mathbf{I}-{\mathbf{P}}_{\tau\left({1}\right)}}\right){\mathbf{P}}_{\tau\left({2}\right)}\cdots{\mathbf{P}}_{\tau\left({t}\right)}}\right)=a_{t}\,,$
showing $(a_{t})_{t}$ is a non-increasing sequence. Thus, for all $k\geq 2$,
$2\mathop{\mathbb{E}}_{\tau}f_{\tau(k)}(\mathbf{w}_{k-1})$ $=\mathop{\mathbb{E}}_{\tau}\left\|{\left({\mathbf{I}-{\mathbf{P}}_{\tau\left({k}\right)}}\right)\mathbf{w}_{k-1}}\right\|^{2}\leq\left\|{\mathbf{w}_{\star}}\right\|^{2}\mathop{\mathbb{E}}_{\tau}a_{k}\leq\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}}{k-1}\sum_{t=2}^{k}\mathop{\mathbb{E}}_{\tau}a_{t}$
$=\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}}{k-1}\sum_{t=2}^{k}\mathop{\mathbb{E}}_{\tau}\left[{\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({t}\right)}\cdots{\mathbf{P}}_{\tau\left({2}\right)}\cdots{\mathbf{P}}_{\tau\left({t}\right)}}\right)-\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({t}\right)}\cdots{\mathbf{P}}_{\tau\left({2}\right)}{\mathbf{P}}_{\tau\left({1}\right)}{\mathbf{P}}_{\tau\left({2}\right)}\cdots{\mathbf{P}}_{\tau\left({t}\right)}}\right)}\right]$
$\left[\begin{subarray}{c}\text{exchangeable}\end{subarray}\right]$ $=\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}}{k-1}\sum_{t=2}^{k}\mathop{\mathbb{E}}_{\tau}\left[{\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({t-1}\right)}\cdots{\mathbf{P}}_{\tau\left({1}\right)}\cdots{\mathbf{P}}_{\tau\left({t-1}\right)}}\right)-\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({t}\right)}\cdots{\mathbf{P}}_{\tau\left({1}\right)}\cdots{\mathbf{P}}_{\tau\left({t}\right)}}\right)}\right]$
$\left[\begin{subarray}{c}\text{telescoping}\end{subarray}\right]$ $=\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}}{k-1}\mathop{\mathbb{E}}_{\tau}\left[{\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({1}\right)}}\right)-\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({k}\right)}\cdots{\mathbf{P}}_{\tau\left({1}\right)}\cdots{\mathbf{P}}_{\tau\left({k}\right)}}\right)}\right]$
$\leq\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}}{k-1}\mathop{\mathbb{E}}_{\tau}\left[{\operatorname{tr}\left({{\mathbf{P}}_{\tau\left({1}\right)}}\right)}\right]=\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}\left({d-\bar{r}}\right)}{k-1}.$
For the second, parameter independent term, note that from Lemma 6, $f_{m}(\cdot)$ is $1$-smooth for all $m\in[T]$, and recall that the iterates $\mathbf{w}_{t}$ follow the SGD dynamics with $\eta=1$ (Reduction 2). Hence, by Lemma 33, without-replacement SGD with $\beta=\eta=1$ guarantees that after $k\geq 1$ gradient steps:
$\mathop{\mathbb{E}}_{\tau}f_{\tau(k)}(\mathbf{w}_{k-1})\leq\frac{e\cdot\left\|{\mathbf{w}_{\star}}\right\|^{2}}{\sqrt[4]{k-1}}.$
Plugging in the (monotonic decreasing) bounds that we just derived in the inequalities from the beginning of this proof, we get
$\mathop{\mathbb{E}}_{\tau}\left[{F_{\tau}\left({k}\right)}\right]$ $\leq 2R^{2}\mathop{\mathbb{E}}_{\tau}f_{\tau(k)}(\mathbf{w}_{k-1})+\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}\leq R^{2}\min\left({\frac{2e\left\|{\mathbf{w}_{\star}}\right\|^{2}}{\sqrt[4]{k-1}},\,\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}\left({d-\bar{r}}\right)}{k-1}}\right)+\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}$
$\leq\min\left({\frac{7}{\sqrt[4]{k-1}},\,\frac{d-\bar{r}+1}{k-1}}\right)\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}\,,$
$\mathbb{E}_{\tau}\left[\mathcal{L}\left(\mathbf{w}_{k}\right)\right]$ $\leq R^{2}\left({\frac{k}{T}2\mathop{\mathbb{E}}_{\tau}f_{\tau(k)}(\mathbf{w}_{k-1})+\frac{T-k}{2T}2\mathop{\mathbb{E}}_{\tau}f_{\tau(k+1)}(\mathbf{w}_{k})}\right)+\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}$
$\leq\left({\frac{k}{T}+\frac{T-k}{2T}}\right)\min\left({\frac{2e}{\sqrt[4]{k-1}},\,\frac{d-\bar{r}}{k-1}}\right)\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}+\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}$
$=\frac{T+k}{2T}\min\left({\frac{2e}{\sqrt[4]{k-1}},\,\frac{d-\bar{r}}{k-1}}\right)\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}+\frac{\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{k}$
$\left[\begin{subarray}{c}k\leq T\end{subarray}\right]$ $\leq\min\left({\frac{7}{\sqrt[4]{k-1}},\,\frac{d-\bar{r}+1}{k-1}}\right)\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}\,.$
Appendix E Proofs of Last-Iterate SGD Bounds (Section 5)
In this section we provide proofs and full technical details of our upper bounds for least squares SGD. We begin by recording a few elementary well-known facts, which can be found in e.g., Bubeck (2015). We provide proof for completeness.
Lemma 30 (Fundamental regret inequality for gradient descent).
Let $\mathbf{w}_{0}\in\mathbb{R}^{d},\eta>0$, and suppose $\mathbf{w}_{t+1}=\mathbf{w}_{t}-\eta\mathbf{g}_{t}$ for all $t$, where $\mathbf{g}_{0},\ldots,\mathbf{g}_{T}\in\mathbb{R}^{d}$ are arbitrary vectors. Then for any $\tilde{\mathbf{w}}\in\mathbb{R}^{d}$ it holds that:
$\sum_{t=0}^{T}\mathbf{g}_{t}^{\top}(\mathbf{w}_{t}-\tilde{\mathbf{w}})\leq\frac{\|\mathbf{w}_{0}-\tilde{\mathbf{w}}\|^{2}}{2\eta}+\frac{\eta}{2}\sum_{t=0}^{T}\|\mathbf{g}_{t}\|^{2}.$
Proof.
Observe,
$\left\|{\mathbf{w}_{t+1}-\tilde{\mathbf{w}}}\right\|^{2}$ $=\left\|{\mathbf{w}_{t}-\tilde{\mathbf{w}}}\right\|^{2}-2\eta\mathbf{g}_{t}^{\top}(\mathbf{w}_{t}-\tilde{\mathbf{w}})+\eta^{2}\left\|{\mathbf{g}_{t}}\right\|^{2}$
$\iff\mathbf{g}_{t}^{\top}(\mathbf{w}_{t}-\tilde{\mathbf{w}})$ $=\frac{1}{2\eta}\left(\left\|{\mathbf{w}_{t}-\tilde{\mathbf{w}}}\right\|^{2}-\left\|{\mathbf{w}_{t+1}-\tilde{\mathbf{w}}}\right\|^{2}\right)+\frac{\eta}{2}\left\|{\mathbf{g}_{t}}\right\|^{2}.$
Summing the above over $t=0,\ldots,T$ and telescoping the sum leads to,
$\sum_{t=0}^{T}\mathbf{g}_{t}^{\top}(\mathbf{w}_{t}-\tilde{\mathbf{w}})$ $=\frac{1}{2\eta}\left(\left\|{\mathbf{w}_{0}-\tilde{\mathbf{w}}}\right\|^{2}-\left\|{\mathbf{w}_{T+1}-\tilde{\mathbf{w}}}\right\|^{2}\right)+\frac{\eta}{2}\sum_{t=0}^{T}\left\|{\mathbf{g}_{t}}\right\|^{2}$
$\leq\frac{\left\|{\mathbf{w}_{0}-\tilde{\mathbf{w}}}\right\|^{2}}{2\eta}+\frac{\eta}{2}\sum_{t=0}^{T}\left\|{\mathbf{g}_{t}}\right\|^{2},$
which completes the proof.
Lemma 31 (Descent lemma).
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be $\beta$-smooth for $\beta>0$, and suppose $\min_{\mathbf{w}}f(\mathbf{w})\in\mathbb{R}$ is attained. Then, for any $\eta>0$, $\mathbf{w}\in\mathbb{R}^{d}$, we have for $\mathbf{w}^{+}=\mathbf{w}-\eta\nabla f(\mathbf{w})$:
$f(\mathbf{w}^{+})\leq f(\mathbf{w})-\eta\left(1-\frac{\eta\beta}{2}\right)\|\nabla f(\mathbf{w})\|^{2}.$
Furthermore, for any $\mathbf{w}_{\star}\in\operatorname*{argmin}_{\mathbf{w}}f(\mathbf{w})$, it holds that:
$\|\nabla f(\mathbf{w})\|^{2}\leq 2\beta\left(f(\mathbf{w})-f(\mathbf{w}_{\star})\right).$
Proof.
Observe, by $\beta$-smoothness:
$f(\mathbf{w}^{+})$ $\leq f(\mathbf{w})+\nabla f(\mathbf{w})\cdot(\mathbf{w}^{+}-\mathbf{w})+\frac{\beta}{2}\|\mathbf{w}^{+}-\mathbf{w}\|^{2}$
$=f(\mathbf{w})-\eta\nabla f(\mathbf{w})\cdot\nabla f(\mathbf{w})+\frac{\beta}{2}\eta^{2}\|\nabla f(\mathbf{w})\|^{2}$
$=f(\mathbf{w})-\eta\left(1-\frac{\eta\beta}{2}\right)\|\nabla f(\mathbf{w})\|^{2},$
which proves the first claim. For the second claim, apply the above inequality with $\eta=1/\beta$, which gives
$f(\mathbf{w}^{+})$ $\leq f(\mathbf{w})-\frac{1}{2\beta}\|\nabla f(\mathbf{w})\|^{2}$
$\iff\|\nabla f(\mathbf{w})\|^{2}$ $\leq 2\beta\left(f(\mathbf{w})-f(\mathbf{w}^{+})\right).$
The second claim now follows by using the fact that $f(\mathbf{w}_{\star})\leq f(\mathbf{w}^{+})$.
E.1 Proofs for Section 5.1
As discussed in the main text, our results hold for a wider range of step sizes compared to the classical SGD bounds in the smooth realizable setting. This is enabled due to the following lemma.
Lemma 32.
Assume that $f(\mathbf{w})=\frac{1}{2}\left\|{\mathbf{A}\mathbf{w}-\mathbf{b}}\right\|^{2}$ for some matrix $\mathbf{A}$ and vector $\mathbf{b}$, and let $\mathbf{w}_{\star}\in\mathbb{R}^{d}$ be such that $f(\mathbf{w}_{\star})=0$. Then, we have:
$2f(\mathbf{w})$ $=\nabla f(\mathbf{w})^{\top}(\mathbf{w}-\mathbf{w}_{\star})\,,$
and for any $\mathbf{z}\in\mathbb{R}^{d}$ and $\gamma>0$:
$(2-\gamma)f(\mathbf{w})-\frac{1}{\gamma}f(\mathbf{z})$ $\leq\nabla f(\mathbf{w})^{\top}(\mathbf{w}-\mathbf{z})\,.$
Proof.
For any $\mathbf{w}\in\mathbb{R}^{d}$, since $\mathbf{A}\mathbf{w}_{\star}=\mathbf{b}$ and $f(\mathbf{w})=\frac{1}{2}\|\mathbf{A}(\mathbf{w}-\mathbf{w}_{\star})\|^{2}$, we have:
$\nabla f(\mathbf{w})^{\top}(\mathbf{w}-\mathbf{z})$ $=\big{\langle}\mathbf{A}^{\top}\mathbf{A}(\mathbf{w}-\mathbf{w}_{\star}),\mathbf{w}-\mathbf{z}\big{\rangle}$
$=\big{\langle}\mathbf{A}^{\top}\mathbf{A}(\mathbf{w}-\mathbf{w}_{\star}),\mathbf{w}-\mathbf{w}_{\star}\big{\rangle}-\big{\langle}\mathbf{A}^{\top}\mathbf{A}(\mathbf{w}-\mathbf{w}_{\star}),\mathbf{z}-\mathbf{w}_{\star}\big{\rangle}$
$=\big{\langle}\mathbf{A}\mathbf{w}-\mathbf{b},\mathbf{A}\mathbf{w}-\mathbf{b}\big{\rangle}-\big{\langle}\mathbf{A}\mathbf{w}-\mathbf{b},\mathbf{A}\mathbf{z}-\mathbf{b}\big{\rangle}$
$=2f(\mathbf{w})-\big{\langle}\mathbf{A}\mathbf{w}-\mathbf{b},\mathbf{A}\mathbf{z}-\mathbf{b}\big{\rangle}.$
Plugging in $\mathbf{z}=\mathbf{w}_{\star}$, the second term vanishes (since $\mathbf{A}\mathbf{w}_{\star}-\mathbf{b}=\mathbf{b}-\mathbf{b}=\mathbf{0}$) and the first claim follows. For the second claim, note that by Young’s inequality:
$\nabla f(\mathbf{w})^{\top}(\mathbf{w}-\mathbf{z})$ $=2f(\mathbf{w})-\big{\langle}\mathbf{A}\mathbf{w}-\mathbf{b},\mathbf{A}\mathbf{z}-\mathbf{b}\big{\rangle}$
$\geq 2f(\mathbf{w})-\frac{\gamma}{2}\left\|{\mathbf{A}\mathbf{w}-\mathbf{b}}\right\|^{2}-\frac{1}{2\gamma}\left\|{\mathbf{A}\mathbf{z}-\mathbf{b}}\right\|^{2}=(2-\gamma)f(\mathbf{w})-\frac{1}{\gamma}f(\mathbf{z})\,.$
  • Recall Lemma 11.Consider the $\beta$-smooth, realizable Setup 1, and let $T\geq 1$, $(i_{0},\ldots,i_{T})\in\mathcal{I}^{T+1}$ be an arbitrary sequence of indices in $\mathcal{I}$, and ${\mathbf{w}_{0}\in\mathbb{R}^{d}}$ be an arbitrary initialization. Then, the gradient descent iterates given by ${\mathbf{w}_{t+1}\leftarrow\mathbf{w}_{t}-\eta\nabla f(\mathbf{w}_{t};i_{t})}$ for a step size $\eta<2/\beta$, hold:
    $\sum_{t=0}^{T}f\left(\mathbf{w}_{t};i_{t}\right)\leq\frac{\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{2\eta(2-\eta\beta)}\,.$
Proof.
Denote $f_{t}(\mathbf{w})\triangleq f(\mathbf{w};i_{t})$, and observe by Lemma 30;
$\sum_{t=0}^{T}\left\langle\nabla f_{t}(\mathbf{w}_{t}),\mathbf{w}_{t}-\mathbf{w}_{\star}\right\rangle$ $\leq\frac{\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{2\eta}+\frac{\eta}{2}\sum_{t=0}^{T}\left\|{\nabla f_{t}(\mathbf{w}_{t})}\right\|^{2}$
$\leq\frac{\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{2\eta}+\eta\beta\sum_{t=0}^{T}f_{t}(\mathbf{w}_{t})-f_{t}(\mathbf{w}_{\star})=\frac{\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{2\eta}+\eta\beta\sum_{t=0}^{T}f_{t}(\mathbf{w}_{t})\,,$
where the second inequality follows from Lemma 31. On the other hand, by Lemma 32,
$\sum_{t=0}^{T}\left\langle\nabla f_{t}(\mathbf{w}_{t}),\mathbf{w}_{t}-\mathbf{w}_{\star}\right\rangle=\sum_{t=0}^{T}2f_{t}(\mathbf{w}_{t})\,.$
Combining the two displays above, it follows that
$(2-\eta\beta)\sum_{t=0}^{T}f_{t}(\mathbf{w}_{t})\leq\frac{\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{2\eta}\,,$
and the result follows after dividing by $(2-\eta\beta)$.
  • Recall Lemma 12.Consider the $\beta$-smooth, realizable Setup 1. Let $T\geq 1$. Assume $\mathcal{P}$ is a distribution over $\mathcal{I}^{T+1}$ such that for every $0\leq t\leq\tau_{1}\leq\tau_{2}\leq T$, the following holds: For any ${i_{0},\ldots i_{t-1}\in\mathcal{I}^{t}},i\in\mathcal{I}$, $\Pr(i_{\tau_{1}}=i|i_{0},\ldots,i_{t-1})=\Pr(i_{\tau_{2}}=i|i_{0},\ldots,i_{t-1})$. Then, for any initialization $\mathbf{w}_{0}\in\mathbb{R}^{d}$, with-replacement SGD (Eq. 2) with step-size $\eta<2/\beta$, holds:
    $\mathbb{E}f(\mathbf{w}_{T},i_{T})\leq(eT)^{\eta\beta\left(1-\eta\beta/4\right)}\mathbb{E}\left[\tfrac{1}{T+1}{\sum}_{t=0}^{T}f(\mathbf{w}_{t};i_{t})\right],$
    where the expectation is taken with respect to $i_{0},\ldots,i_{T}$ sampled from $\mathcal{P}$.
Proof.
Denote $f_{t}(\mathbf{w})\!\triangleq\!f(\mathbf{w};i_{t}),\,\mathbf{g}_{t}\!\triangleq\!\nabla f_{t}(\mathbf{w}_{t})$, and observe that by Lemma 30, $\forall\mathbf{z}\in\mathbb{R}^{d},\,t\leq T$ (w.p. $1$):
$\sum_{t=T-k}^{T}\left\langle\mathbf{g}_{t},\mathbf{w}_{t}-\mathbf{z}\right\rangle$ $\leq\frac{\left\|{\mathbf{w}_{T-k}-\mathbf{z}}\right\|^{2}}{2\eta}+\frac{\eta}{2}\sum_{t=T-k}^{T}\left\|{\mathbf{g}_{t}}\right\|^{2}$
[…]
$=\frac{\left\|{\mathbf{w}_{T-k}-\mathbf{z}}\right\|^{2}}{2\eta}+\eta\beta\sum_{t=T-k}^{T}f_{t}(\mathbf{w}_{t})-f_{t}(\mathbf{z})+f_{t}(\mathbf{z})-f_{t}(\mathbf{w}_{\star})\,.$
By Lemma 32, this implies for any $\gamma>0$:
$\sum_{t=T-k}^{T}(2-\gamma-\eta\beta)f_{t}(\mathbf{w}_{t})-\left(\frac{1}{\gamma}-\eta\beta\right)f_{t}(\mathbf{z})$ $=\sum_{t=T-k}^{T}\!\Big{(}{(2-\gamma)f_{t}(\mathbf{w}_{t})-\frac{1}{\gamma}f_{t}(\mathbf{z})}\Big{)}+\eta\beta\!\sum_{t=T-k}^{T}f_{t}(\mathbf{z})-f_{t}(\mathbf{w}_{t})$
[Lemma 32]delimited-[]Lemma $\end{subarray}\right]$ $\leq\sum_{t=T-k}^{T}\left\langle\mathbf{g}_{t},\mathbf{w}_{t}-\mathbf{z}\right\rangle+\eta\beta\sum_{t=T-k}^{T}f_{t}(\mathbf{z})-f_{t}(\mathbf{w}_{t})$
$\left[\begin{subarray}{c}\text{above}\end{subarray}\right]$ $\leq\frac{\left\|{\mathbf{w}_{T-k}-\mathbf{z}}\right\|^{2}}{2\eta}+\eta\beta\sum_{t=T-k}^{T}f_{t}(\mathbf{z})-\underbrace{f_{t}(\mathbf{w}_{\star})}_{=0}$
$\implies(2-\gamma-\eta\beta)\sum_{t=T-k}^{T}f_{t}(\mathbf{w}_{t})$ $\leq\frac{\left\|{\mathbf{w}_{T-k}-\mathbf{z}}\right\|^{2}}{2\eta}+\frac{1}{\gamma}\sum_{t=T-k}^{T}f_{t}(\mathbf{z})\,.$
Now, set $\mathbf{z}=\mathbf{w}_{T-k}$ and take expectations to obtain:
$(2-\gamma-\eta\beta)\sum_{t=T-k}^{T}\mathbb{E}f_{t}(\mathbf{w}_{t})$ $\leq 0+\frac{1}{\gamma}\sum_{t=T-k}^{T}\mathbb{E}f_{t}(\mathbf{w}_{T-k})$
$\frac{1}{k+1}\sum_{t=T-k}^{T}\mathbb{E}f_{t}(\mathbf{w}_{t})$ $\leq\frac{1}{\big{(}{k+1}\big{)}\gamma(2-\gamma-\eta\beta)}\sum_{t=T-k}^{T}\mathbb{E}f_{t}(\mathbf{w}_{T-k})\,.$
Defining $S_{k}\triangleq\frac{1}{k+1}\sum_{t=T-k}^{T}f_{t}(\mathbf{w}_{t})$, implies that
$(k+1)S_{k}-kS_{k-1}$ $=\sum_{t=T-k}^{T}f_{t}(\mathbf{w}_{t})-\sum_{t=T-k+1}^{T}f_{t}(\mathbf{w}_{t})=f_{T-k}(\mathbf{w}_{T-k})\,,$
and by the assumption on the distribution $\mathcal{P}$ it follows that $\mathbb{E}f_{T-k}(\mathbf{w}_{T-k})=\mathbb{E}f_{t}(\mathbf{w}_{T-k})$ for any $t\geq T-k$. Thus, combined with our previous display,
$\mathbb{E}S_{k}$ $\leq\frac{1}{\big{(}{k+1}\big{)}\gamma(2-\gamma-\eta\beta)}\sum_{t=T-k}^{T}\mathbb{E}f_{t}(\mathbf{w}_{T-k})=\frac{1}{\big{(}{k+1}\big{)}\gamma(2-\gamma-\eta\beta)}\sum_{t=T-k}^{T}\Big{(}{(k+1)\mathbb{E}S_{k}-k\mathbb{E}S_{k-1}}\Big{)}$
$=\frac{1}{\gamma(2-\gamma-\eta\beta)}\Big{(}{(k+1)\mathbb{E}S_{k}-k\mathbb{E}S_{k-1}}\Big{)}\,.$
Rearranging, denoting $c\triangleq\gamma(2-\gamma-\eta\beta)$, and requiring $c\in(0,1)$, we get
$\frac{k}{c}\mathbb{E}S_{k-1}$ $\leq\left({\frac{k+1}{c}-1}\right)\mathbb{E}S_{k}$
$\iff\mathbb{E}S_{k-1}$ $\leq\frac{k+1-c}{k}\mathbb{E}S_{k}$
$\implies\mathbb{E}f_{T}(\mathbf{w}_{T})=\mathbb{E}S_{0}$ $\leq\prod_{k=1}^{T}\left({1+\frac{1-c}{k}}\right)\mathbb{E}S_{T}$
$\left[\begin{subarray}{c}1+x\leq e^{x},\forall x\geq 0\end{subarray}\right]$ $\leq\exp\left({\sum_{k=1}^{T}\frac{1-c}{k}}\right)\mathbb{E}S_{T}$
$=\exp\left({\left({1-c}\right)\sum_{k=1}^{T}\frac{1}{k}}\right)\cdot\mathbb{E}S_{T}\leq\exp\big{(}{\left({1-c}\right)\left({1+\log T}\right)}\big{)}\mathbb{E}S_{T}$
$=\left({eT}\right)^{1-c}\cdot\mathbb{E}\left[\frac{1}{T+1}\sum_{t=0}^{T}f_{t}(\mathbf{w}_{t})\right]\,.$ (7)
Now, getting the “best” rate requires maximizing $c=\gamma(2-\gamma-\eta\beta)$. To this end, we choose $\gamma=1-\frac{\eta\beta}{2}$, which implies $c=\big{(}{1-\frac{\eta\beta}{2}}\big{)}^{2}$ (under the $\eta<\frac{2}{\beta}$ condition, we now have both $\gamma>0$ and $c\in(0,1)$ as required above). Then, $1-c=\eta\beta\big{(}{1-\frac{\eta\beta}{4}}\big{)}$, and we finally get the required
$\mathbb{E}f_{T}(\mathbf{w}_{T})$ $=\left({eT}\right)^{\eta\beta\big{(}{1-\frac{\eta\beta}{4}}\big{)}}\cdot\frac{1}{T+1}\sum_{t=0}^{T}f_{t}(\mathbf{w}_{t})\,.$
E.2 Proving Theorem 13
  • Recall Theorem 13.Consider the $\beta$-smooth, realizable Setup 1. Define for all $T\!\geq\!2$, $\hat{f}_{0:T}(\mathbf{w})\!\triangleq\!\frac{1}{T+1}\sum_{t=0}^{T}f(\mathbf{w};\pi_{t})$. Then, without-replacement SGD (Eq. (3)) with step-size $\eta<{2}/{\beta}$, holds:
    $\mathbb{E}_{\pi}\hat{f}_{0:T}(\mathbf{w}_{T})\leq\frac{eD^{2}}{\eta(2-\eta\beta)T^{1-\eta\beta\left(1-\eta\beta/4\right)}}+\frac{4\beta^{2}\eta D^{2}}{T}\,,\quad\forall T=2,\dots,n-1\,,$
    where $D\triangleq\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|$. In particular,
    $\eta=\frac{1}{\beta\log T}\Longrightarrow\mathbb{E}\hat{f}_{0:T-1}(\mathbf{w}_{T})\leq{\displaystyle\frac{14\beta D^{2}\log T}{T}}\,,\quad\quad\eta=\frac{1}{\beta}\Longrightarrow\mathbb{E}\hat{f}_{0:T-1}(\mathbf{w}_{T})\leq\frac{7\beta D^{2}}{\sqrt[4]{T}}\,.$
The proof of our theorem follows by a combination of two lemmas. The first, stated below, establishes a bound on the expected “next sample” loss and follows immediately by combining Lemmas 12 and 11 (notice that $\eta<\frac{2}{\beta}\Longrightarrow\exp\big{(}{\eta\beta\big{(}{1-\frac{\eta\beta}{4}}\big{)}}\big{)}\mapsto\exp\big{(}{z\big{(}{1-\frac{z}{4}}\big{)}}\big{)}$ for $z\in(0,2)$, which is monotonic increasing and upper bounded by $e$).
Lemma 33.
For any step-size $\eta<2/\beta$ and initialization $\mathbf{w}_{0}\in\mathbb{R}^{d}$, without-replacement SGD Eq. 3 satisfies, for all $11\leq T\leq n-$ 1:
$\mathbb{E}_{\pi}f(\mathbf{w}_{T};\pi_{T})$ $\leq e^{\eta\beta\left(1-\frac{\eta\beta}{4}\right)}T^{\eta\beta\left(1-\frac{\eta\beta}{4}\right)}\mathbb{E}_{\pi}\left[\frac{1}{T+1}\sum_{t=0}^{T}f(\mathbf{w}_{t};\pi_{t})\right]\leq\frac{e\cdot\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{2\eta(2-\eta\beta)T^{1-\eta\beta\left(1-\frac{\eta\beta}{4}\right)}}\,.$
Next, we consider the “empirical loss” objective. Given any permutation $\pi\in\mathcal{I}\leftrightarrow\mathcal{I}$, define:
$\hat{f}_{0:t}(\mathbf{w})\triangleq\frac{1}{t+1}\sum_{i=0}^{t}f(\mathbf{w};\pi_{i}).$
In the without-replacement setup, our optimization objective is the expected empirical loss $\mathbb{E}_{\pi}\hat{f}_{0:t}(\mathbf{w})$, which, when $t=n$, satisfies $\mathbb{E}_{\pi}\hat{f}_{0:t}(\mathbf{w})=\mathbb{E}_{\pi}\bar{f}(\mathbf{w})$. Our second lemma (given next) bounds the expected empirical loss w.r.t. the next sample loss. This is the crux of extending our with-replacement upper bound to the without-replacement setup.
Lemma 34.
For without-replacement SGD Eq. 3 with step size $\eta\leq 2/\beta$, for all $1\leq T\leq n$, we have that the following holds:
$\mathbb{E}_{\pi}\hat{f}_{0:T}(\mathbf{w}_{T})$ $\leq 2\mathbb{E}_{\pi}f(\mathbf{w}_{T};\pi_{T})+\frac{4\beta^{2}\eta\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{T+1}.$
The proof of Lemma 34 builds on an algorithmic stability argument similar to that given in Lei and Ying (2020), combined with the without-replacement stability framework proposed by Sherman et al. (2021). Before turning to the proof given in the next subsection, we quickly prove Theorem 13.
Proof (of Theorem 13).
By Lemmas 34 and 33,
$\mathbb{E}_{\pi}\hat{f}_{0:T}(\mathbf{w}_{T})$ $\leq 2\mathbb{E}_{\pi}f(\mathbf{w}_{T};\pi_{T})+\frac{4\beta^{2}\eta\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{T+1}\leq\frac{e\cdot\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{\eta(2-\eta\beta)}T^{\eta\beta\left(1-\frac{\eta\beta}{4}\right)-1}+\frac{4\beta^{2}\eta\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{T+1}.$
The result for $\eta=\frac{1}{\beta}$ is straightforward. To see the result for $\eta=\frac{1}{\beta\log T}$, notice that in this case,
$\frac{eD^{2}T^{\eta\beta\left(1-\eta\beta/4\right)-1}}{\eta(2-\eta\beta)}=\frac{e\beta D^{2}\log T}{T(2-\frac{1}{\log T})}T^{\frac{1}{\log T}\left(1-\frac{1}{4\log T}\right)}=\frac{\beta D^{2}\log T}{T}\frac{\exp\left(2-\frac{1}{4\log T}\right)}{2-\frac{1}{\log T}}\leq\frac{10\beta D^{2}\log T}{T}.~{}~{}~{}~{}~{}~{}~{}$
E.2.1 Proving Lemma 34
Notation.
We first add a few definitions central to our analysis. Given a permutation $\pi\in\mathcal{I}\leftrightarrow\mathcal{I}$, denote:
[…]
$\mathbf{w}_{\tau}^{\pi}$ $\triangleq\text{The iterate of SGD on step }\tau\text{ when run on permutation }\pi.$
Most commonly, we will use the following special case of the above:
$\mathbf{w}_{\tau}^{\pi(i\leftrightarrow t)}$ $\triangleq\text{The iterate of SGD on step }\tau\text{ when run on }\pi(i\leftrightarrow t).$
When clear from context, we omit $\pi$ from the superscript and simply write $\mathbf{w}_{\tau}^{(i\leftrightarrow t)}$. Concretely, these definitions imply $\mathbf{w}_{0}^{(i\leftrightarrow t)}\triangleq\mathbf{w}_{0}$, and $\forall i,t,\tau\in\mathcal{I}$,
$\mathbf{w}_{\tau+1}^{(i\leftrightarrow t)}$ $=\mathbf{w}_{\tau}^{(i\leftrightarrow t)}-\eta\nabla f\Big{(}{\mathbf{w}_{\tau}^{(i\leftrightarrow t)};\pi(i\leftrightarrow t)_{\tau}}\Big{)}\,.$
We have the following important relation, to be used later in the proof.
Lemma 35.
For all $i,t,\tau\in\mathcal{I},i\leq\tau\leq t$, we have:
$\mathbb{E}_{\pi}f(\mathbf{w}_{\tau};\pi_{i})=\mathbb{E}_{\pi}f(\mathbf{w}_{\tau}^{(i\leftrightarrow t)};\pi(i\leftrightarrow t)_{i})\,.$
Proof.
The proof follows from observing that the random variables $f(\mathbf{w}_{\tau};\pi_{i})$ and $f(\mathbf{w}_{\tau}^{(i\leftrightarrow t)};\pi(i\leftrightarrow t)_{i})$ are distributed identically (the indices $\pi_{i},\pi_{t}$ are exhangeable). Formally, let $\Pi(\mathcal{I})\triangleq\left\{\pi\in\mathcal{I}\leftrightarrow\mathcal{I}\right\}$ be the set of all permutations over $\mathcal{I}$, and observe
$\mathbb{E}_{\pi}f(\mathbf{w}_{\tau}^{(i\leftrightarrow t)};\pi(i\leftrightarrow t)_{i})=\frac{1}{|\Pi(\mathcal{I})|}\sum_{\pi\in\Pi(\mathcal{I})}f(\mathbf{w}_{\tau}^{\pi(i\leftrightarrow t)};\pi(i\leftrightarrow t)_{i})\,.$
On the other hand,
$\mathbb{E}_{\pi}f(\mathbf{w}_{\tau},\pi_{i})=\frac{1}{|\Pi(\mathcal{I})|}\sum_{\pi\in\Pi(\mathcal{I})}f(\mathbf{w}_{\tau}^{\pi};\pi_{i})\,.$
Hence, since there is a one-to-one correspondence between $\pi$ and $\pi(\tau\leftrightarrow i)$, in particular,
$\left\{\pi\mid\pi\in\Pi(\mathcal{I})\right\}=\left\{\pi(i\leftrightarrow t)\mid\pi\in\Pi(\mathcal{I})\right\},$
the result follows.
Our next lemma, originally given in Sherman et al. (2021, Lemma 2 therein), can be thought of as a without-replacement version of the well known stability $\iff$ generalization argument of the with-replacement sampling case (Shalev-Shwartz et al., 2010; Hardt et al., 2016).
Lemma 36.
The iterates of without-replacement SGD Eq. 3, satisfy for all $t$:
$\mathbb{E}_{\pi}\left[f(\mathbf{w}_{t};\pi_{t})-\hat{f}_{0:t-1}(\mathbf{w}_{t})\right]=\frac{1}{t}\sum_{i=0}^{t-1}\mathbb{E}_{\pi}\left[f(\mathbf{w}_{t};\pi_{t})-f(\mathbf{w}_{t}^{(i\leftrightarrow t)};\pi_{t})\right]$
Proof.
We have, by definition of $\hat{f}_{0:t-1}$ and Lemma 35:
$\mathbb{E}_{\pi}\left[\hat{f}_{0:t-1}(\mathbf{w}_{t})\right]=\frac{1}{t}\sum_{i=0}^{t-1}\mathbb{E}_{\pi}\left[f(\mathbf{w}_{t};\pi_{i})\right]=\frac{1}{t}\sum_{i=0}^{t-1}\mathbb{E}_{\pi}\left[f(\mathbf{w}_{t}^{(i\leftrightarrow t)};\pi(i\leftrightarrow t)_{i})\right]=\frac{1}{t}\sum_{i=0}^{t-1}\mathbb{E}_{\pi}\left[f(\mathbf{w}_{t}^{(i\leftrightarrow t)};\pi_{t})\right],$
where the last equality is immediate since by definition, $\pi(i\leftrightarrow t)_{i}=\pi_{t}$. The claim now follows by linearity of expectation.
We are now ready to prove our main lemma. We note that the proof shares some features with that of the with-replacement case (Lemma 38).
Proof (of Lemma 34).
We prove the theorem for every $t$. Any $\beta$-smooth realizable function $h:\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ holds that
$\left|h(\tilde{\mathbf{w}})-h(\mathbf{w})\right|$ $\leq\left|\nabla h(\mathbf{w})^{\top}(\tilde{\mathbf{w}}-\mathbf{w})\right|+\frac{\beta}{2}\left\|{\tilde{\mathbf{w}}-\mathbf{w}}\right\|^{2}$
$\left[\begin{subarray}{c}\text{Young's ineq.}\end{subarray}\right]$ $\leq\frac{1}{2\beta}\left\|{\nabla h(\mathbf{w})}\right\|^{2}+\frac{\beta}{2}\left\|{\tilde{\mathbf{w}}-\mathbf{w}}\right\|^{2}+\frac{\beta}{2}\left\|{\tilde{\mathbf{w}}-\mathbf{w}}\right\|^{2}$
$\leq h(\mathbf{w})+\beta\left\|{\tilde{\mathbf{w}}-\mathbf{w}}\right\|^{2}\,.$ (8)
Hence, by Lemma 36,
$\left|\mathbb{E}_{\pi}\left[f(\mathbf{w}_{t};\pi_{t})-\hat{f}_{0:t-1}(\mathbf{w}_{t})\right]\right|$ $=\left|\frac{1}{t}\sum_{i=0}^{t-1}\mathbb{E}_{\pi}\left[f(\mathbf{w}_{t};\pi_{t})-f(\mathbf{w}_{t}^{(i\leftrightarrow t)};\pi_{t})\right]\right|$
$\left[\begin{subarray}{c}\text{Jensen}\end{subarray}\right]$ $\leq\frac{1}{t}\sum_{i=0}^{t-1}\mathbb{E}_{\pi}\left|{f(\mathbf{w}_{t};\pi_{t})-f(\mathbf{w}_{t}^{(i\leftrightarrow t)};\pi_{t})}\right|$
[…]
$=\mathbb{E}_{\pi}f(\mathbf{w}_{t},\pi_{t})+\frac{\beta}{t}\sum_{i=0}^{t-1}\mathbb{E}_{\pi}\left\|{\mathbf{w}_{t}^{(i\leftrightarrow t)}-\mathbf{w}_{t}}\right\|^{2}.$ (9)
Next, we bound $\left\|{\mathbf{w}_{t}^{(i\leftrightarrow t)}-\mathbf{w}_{t}}\right\|^{2}$. For any $10\leq\tau\leq t-$ 1, we denote $f_{\tau}\triangleq f(\cdot;\pi_{\tau})$, $f_{\tau}^{(i\leftrightarrow t)}\triangleq f(\cdot;\pi(i\leftrightarrow t)_{\tau})$. Observe that for any $\tau$ such that $\tau\neq i$, we have $f_{\tau}=f_{\tau}^{(i\leftrightarrow t)}$, thus, by the non-expansiveness of gradient steps in the convex and $\beta$-smooth regime when $\eta\leq 2/\beta$ (see Lemma 3.6 in Hardt et al., 2016):
$\tau\leq i$ $\implies\left\|{\mathbf{w}_{\tau}^{(i\leftrightarrow t)}-\mathbf{w}_{\tau}}\right\|=0,$
$i<\tau$ $\implies\left\|{\mathbf{w}_{\tau+1}^{(i\leftrightarrow t)}-\mathbf{w}_{\tau+1}}\right\|^{2}\leq\left\|{\mathbf{w}_{i+1}^{(i\leftrightarrow t)}-\mathbf{w}_{i+1}}\right\|^{2}.$
Further,
$\left\|{\mathbf{w}_{i+1}^{(i\leftrightarrow t)}-\mathbf{w}_{i+1}}\right\|^{2}$ $=\left\|{\mathbf{w}_{i}^{(i\leftrightarrow t)}-\eta\nabla f_{i}^{(i\leftrightarrow t)}(\mathbf{w}_{i}^{(i\leftrightarrow t)})-\left({\mathbf{w}_{i}-\eta\nabla f_{i}(\mathbf{w}_{i})}\right)}\right\|^{2}$
$\left[\begin{subarray}{c}\mathbf{w}_{i}^{(i\leftrightarrow t)}=\mathbf{w}_{i}\end{subarray}\right]$ $=\eta^{2}\left\|{\nabla f_{i}^{(i\leftrightarrow t)}(\mathbf{w}_{i}^{(i\leftrightarrow t)})-\nabla f_{i}(\mathbf{w}_{i})}\right\|^{2}$
$\left[\begin{subarray}{c}\text{Jensen}\end{subarray}\right]$ $\leq 2\eta^{2}\left\|{\nabla f_{i}^{(i\leftrightarrow t)}(\mathbf{w}_{i}^{(i\leftrightarrow t)})}\right\|^{2}+2\eta^{2}\left\|{\nabla f_{i}(\mathbf{w}_{i})}\right\|^{2}$
$\leq 4\beta\eta^{2}f_{i}^{(i\leftrightarrow t)}(\mathbf{w}_{i}^{(i\leftrightarrow t)})+4\beta\eta^{2}f_{i}(\mathbf{w}_{i})\,,$
and by Lemma 35 $\mathbb{E}f_{i}(\mathbf{w}_{i})=\mathbb{E}f_{i}^{(i\leftrightarrow t)}(\mathbf{w}_{i}^{(i\leftrightarrow t)})$. Hence,
$\mathbb{E}\left\|{\mathbf{w}_{t}^{(i\leftrightarrow t)}-\mathbf{w}_{t}}\right\|^{2}\leq\mathbb{E}\left\|{\mathbf{w}_{i+1}^{(i\leftrightarrow t)}-\mathbf{w}_{i+1}}\right\|^{2}$ $\leq 8\beta\eta^{2}\mathbb{E}f_{i}(\mathbf{w}_{i})\,.$
Now,
$\frac{\beta}{t}\sum_{i=0}^{t-1}\mathbb{E}_{\pi}\left\|{\mathbf{w}_{t}^{(i\leftrightarrow t)}-\mathbf{w}_{t}}\right\|^{2}$ $\leq\left(8\beta^{2}\eta^{2}\right)\mathbb{E}\left[\frac{1}{t}\sum_{i=0}^{t-1}f_{i}(\mathbf{w}_{i})\right],$
which, when combined with Proof yields:
$\left|\mathbb{E}_{\pi}\left[f(\mathbf{w}_{t};\pi_{t})-\hat{f}_{0:t-1}(\mathbf{w}_{t})\right]\right|$ $\leq\mathbb{E}_{\pi}f(\mathbf{w}_{t};\pi_{t})+\left(8\beta^{2}\eta^{2}\right)\mathbb{E}\left[\frac{1}{t}\sum_{i=0}^{t-1}f_{i}(\mathbf{w}_{i})\right].$
Finally, by the regret bound given in Lemma 11, $\sum_{i=0}^{t-1}f_{i}(\mathbf{w}_{i})\leq\frac{\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{2\eta(2-\eta\beta)}$, and therefore,
$\left|\mathbb{E}_{\pi}\left[f(\mathbf{w}_{t};\pi_{t})-\hat{f}_{0:t-1}(\mathbf{w}_{t})\right]\right|$ $\leq\mathbb{E}_{\pi}f(\mathbf{w}_{t};\pi_{t})+\frac{4\beta^{2}\eta\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{(2-\eta\beta)t}$
$\implies\mathbb{E}\hat{f}_{0:t-1}(\mathbf{w}_{t})$ $\leq 2\mathbb{E}_{\pi}f(\mathbf{w}_{t};\pi_{t})+\frac{4\beta^{2}\eta\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{(2-\eta\beta)t}.$
Finally, since $\hat{f}_{0:t}=\frac{t}{t+1}\hat{f}_{0:t-1}+\frac{1}{t+1}f_{t}$, we obtain
$\mathbb{E}\hat{f}_{0:t}(\mathbf{w}_{t})=\frac{t}{t+1}\mathbb{E}\hat{f}_{0:t-1}(\mathbf{w}_{t})+\frac{1}{t+1}\mathbb{E}f_{t}(\mathbf{w}_{t})$ $\leq\frac{2t+1}{t+1}\mathbb{E}_{\pi}f(\mathbf{w}_{t};\pi_{t})+\frac{4\beta^{2}\eta\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{(2-\eta\beta)(t+1)},$
which completes the proof.
Appendix F Proofs of Extensions (Section 6)
  • Recall Reduction 3.Consider $T$ arbitrary (nonempty) closed convex sets $\mathcal{C}_{1},\dots,\mathcal{C}_{T}$, initial point $\mathbf{w}_{0}\in\mathbb{R}^{d}$, and ordering $\tau$. Define $f_{m}(\mathbf{w})=\frac{1}{2}\left\|{\mathbf{w}-{\Pi}_{m}(\mathbf{w})}\right\|^{2},\forall m\in\left[{T}\right]$. Then,
    1. (i)
      $f_{m}$ is convex and $1$-smooth.
    2. (ii)
      The POCS update is equivalent to an SGD step: $\mathbf{w}_{t}={\Pi}_{\tau(t)}(\mathbf{w}_{t-1})=\mathbf{w}_{t-1}-\nabla_{\mathbf{w}}f_{\tau(t)}(\mathbf{w}_{t-1})$.
Proof.
First, by Theorem 1.5.5 in Facchinei and Pang (2003), $f_{m}$ is continuously differentiable and for every $\mathbf{w}\in\mathbb{R}^{d},m\in\left[{T}\right]$, $\nabla f_{m}(\mathbf{w})=\mathbf{w}-{\Pi}_{m}(\mathbf{w})$. Plugging in $\nabla f_{\tau(t)}(\mathbf{w}_{t-1})$ into an appropriate SGD step, we get
$\mathbf{w}_{t}=\mathbf{w}_{t-1}-\nabla_{\mathbf{w}}f_{\tau(t)}(\mathbf{w}_{t-1})=\mathbf{w}_{t-1}-\left(\mathbf{w}_{t-1}-{\Pi}_{\tau(t)}(\mathbf{w}_{t-1})\right)={\Pi}_{\tau(t)}(\mathbf{w}_{t-1})\,,$
and the second part of the lemma follows. In addition, $\forall\mathbf{x},\mathbf{w}\in\mathbb{R}^{d}$, we prove convexity by using a projection inequality (also from Theorem 1.5.5 in Facchinei and Pang, 2003). That is,
$f_{m}(\mathbf{x})-f_{m}(\mathbf{w})-\langle\nabla f_{m}(\mathbf{w}),\mathbf{x}-\mathbf{w}\rangle=\frac{1}{2}\|\mathbf{x}-{\Pi}_{m}(\mathbf{x})\|^{2}-\frac{1}{2}\|\mathbf{w}-{\Pi}_{m}(\mathbf{w})\|^{2}-\langle\mathbf{w}-{\Pi}_{m}(\mathbf{w}),\mathbf{x}-\mathbf{w}\rangle$
$\hskip 41.00012pt=\frac{1}{2}\|\mathbf{x}-{\Pi}_{m}(\mathbf{x})\|^{2}-\frac{1}{2}\|\mathbf{w}-{\Pi}_{m}(\mathbf{w})\|^{2}-\langle\mathbf{w}-{\Pi}_{m}(\mathbf{w}),\mathbf{x}-{\Pi}_{m}(\mathbf{x})\rangle$
$\hskip 75.0001pt+\langle\mathbf{w}-{\Pi}_{m}(\mathbf{w}),{\Pi}_{m}(\mathbf{w})-{\Pi}_{m}(\mathbf{x})\rangle+\langle\mathbf{w}-{\Pi}_{m}(\mathbf{w}),\mathbf{w}-{\Pi}_{m}(\mathbf{w})\rangle$
$\hskip 41.00012pt\geq\frac{1}{2}\|\mathbf{x}-{\Pi}_{m}(\mathbf{x})\|^{2}-\frac{1}{2}\|\mathbf{w}-{\Pi}_{m}(\mathbf{w})\|^{2}-\langle\mathbf{w}-{\Pi}_{m}(\mathbf{w}),\mathbf{x}-{\Pi}_{m}(\mathbf{x})\rangle+0+\|\mathbf{w}-{\Pi}_{m}(\mathbf{w})\|^{2}$
$\hskip 41.00012pt=\frac{1}{2}\|\mathbf{x}-{\Pi}_{m}(\mathbf{x})-\mathbf{w}+{\Pi}_{m}(\mathbf{w})\|^{2}\geq 0\,.$
For the $1$-smoothness,
$\left\|{\nabla f_{m}(\mathbf{x})-\nabla f_{m}(\mathbf{w})}\right\|=\left\|{\mathbf{x}-{\Pi}_{m}(\mathbf{x})-\left({\mathbf{w}-{\Pi}_{m}(\mathbf{w})}\right)}\right\|=\left\|{(\mathbf{I}-{\Pi}_{m})(\mathbf{x})-(\mathbf{I}-{\Pi}_{m})(\mathbf{w})}\right\|\leq\left\|{\mathbf{x}-\mathbf{w}}\right\|\,,$
where we used the non-expansiveness of $\mathbf{I}-{\Pi}_{m}$ (Propositions 4.2, 4.8 in Bauschke et al., 2011).
Lemma 37.
Let $\mathcal{K}\subseteq\mathbb{R}^{d}$ be a nonempty closed and convex set, and $f(\mathbf{w})=\frac{1}{2}\left\|{\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w})}\right\|^{2}$. Then, we have for any $\mathbf{z}\in\mathbb{R}^{d}$ and $\gamma>0$
$(2-\gamma)f(\mathbf{w})-\frac{1}{\gamma}f(\mathbf{z})$ $\leq\nabla f(\mathbf{w})^{\top}(\mathbf{w}-\mathbf{z})\,.$
In addition, for any $\mathbf{u}\in\mathcal{K}$ we have
$2f(\mathbf{w})$ $\leq\nabla f(\mathbf{w})^{\top}(\mathbf{w}-\mathbf{u})\,.$
Proof.
We have $\nabla f(\mathbf{w})=\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w})$. Hence, by Theorem 1.5.5 in Facchinei and Pang (2003),
$\left\langle\nabla f(\mathbf{w}),\mathbf{w}-\mathbf{z}\right\rangle$ $=\left\langle\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w}),\mathbf{w}-\mathbf{z}\right\rangle=\left\langle\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w}),\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w})\right\rangle+\left\langle\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w}),{\Pi}_{\mathcal{K}}(\mathbf{w})-\mathbf{z}\right\rangle$
$=2f(\mathbf{w})+\left\langle\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w}),{\Pi}_{\mathcal{K}}(\mathbf{w})-{\Pi}_{\mathcal{K}}(\mathbf{z})\right\rangle-\left\langle\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w}),\mathbf{z}-{\Pi}_{\mathcal{K}}(\mathbf{z})\right\rangle$
$\geq 2f(\mathbf{w})-\left\langle\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w}),\mathbf{z}-{\Pi}_{\mathcal{K}}(\mathbf{z})\right\rangle\,.$
Plugging in $\mathbf{z}=\mathbf{u}$, the second term vanishes (since $\mathbf{u}-{\Pi}_{\mathcal{K}}(\mathbf{u})=\mathbf{0}$) and the second claim follows.
For the first claim, note that by Young’s inequality:
$\left\langle\nabla f(\mathbf{w}),\mathbf{w}-\mathbf{z}\right\rangle$ $=2f(\mathbf{w})-\left\langle\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w}),\mathbf{z}-{\Pi}_{\mathcal{K}}(\mathbf{z})\right\rangle$
$\geq 2f(\mathbf{w})-\frac{\gamma}{2}\left\|{\mathbf{w}-{\Pi}_{\mathcal{K}}(\mathbf{w})}\right\|^{2}-\frac{1}{2\gamma}\left\|{\mathbf{z}-{\Pi}_{\mathcal{K}}(\mathbf{z})}\right\|^{2}=2f(\mathbf{w})-\gamma f(\mathbf{w})-\frac{1}{\gamma}f(\mathbf{z})\,.$
  • Recall Theorem 14.Consider the same conditions of Reduction 3 and assume a nonempty set intersection $\mathcal{C}_{\star}=\bigcap_{m=1}^{T}\mathcal{C}_{m}\neq\varnothing$. Then, under a random ordering with or without replacement, the expected “residual” of Scheme 4 after $\forall k\geq 1$ iterations (without replacement: $k\in\left[{T}\right]$) is bounded as,
    $\mathbb{E}_{\tau}\Big{[}\frac{1}{2T}{\sum}_{m=1}^{T}\left\|{\mathbf{w}_{k}-{\Pi}_{m}(\mathbf{w}_{k})}\right\|^{2}\Big{]}=\mathbb{E}_{\tau}\Big{[}\frac{1}{2T}{\sum}_{m=1}^{T}\mathrm{dist}^{2}(\mathbf{w}_{k},\mathcal{C}_{m})\Big{]}\leq\frac{7}{\sqrt[4]{k}}\min_{\mathbf{w}\in\mathcal{C}_{\star}}\left\|{\mathbf{w}_{0}-\mathbf{w}}\right\|^{2}\,.$
    Proof.
    The proof largely follows the same steps of Theorems 8 and 9. Let $\tau$ be any random ordering, $\mathbf{w}_{0}\in\mathbb{R}^{d}$ an initialization, and $\mathbf{w}_{1},\ldots,\mathbf{w}_{k}$ be the corresponding iterates produced by Scheme 4. By Reduction 3, these are exactly the (stochastic) gradient descent iterates produced when initializing at $\mathbf{w}_{0}$ and using a step size of $\eta=1$, on the $1$-smooth loss sequence $f_{\tau(1)},\ldots,f_{\tau(k)}$ defined by:
    $f_{m}(\mathbf{w})\triangleq\frac{1}{2}\left\|{\mathbf{w}-{\Pi}_{m}(\mathbf{w}))}\right\|^{2}.$
    Proceeding, we denote the objective function:
    $\bar{f}(\mathbf{w})\triangleq\mathbb{E}_{m\sim\textrm{Unif}([T])}f_{m}(\mathbf{w})=\frac{1}{2T}\sum_{m=1}^{T}\left\|{\mathbf{w}-{\Pi}_{m}(\mathbf{w})}\right\|^{2}\,.$
    Now, for a with-replacement ordering $\tau$, invoke Theorem 10, except we use Lemma 37 in the proof instead of Lemma 32, to obtain:
    $\mathbb{E}_{\tau}\bar{f}(\mathbf{w}_{k})\leq\frac{e}{2\sqrt[4]{k}}\min_{\mathbf{w}\in\mathcal{C}_{\star}}\left\|{\mathbf{w}_{0}-\mathbf{w}}\right\|^{2}\,,$ ($\tau$ with-replacement)
    which completes the proof for the with-replacement case.
For a without-replacement ordering $\tau$, invoke Theorem 13 (with $\eta=1/\beta$), except again we use Lemma 37 in the proof instead of Lemma 32, to obtain:
$\mathbb{E}_{\tau}\hat{f}_{0:k-1}(\mathbf{w}_{k})\triangleq\mathbb{E}_{\tau}\Big{[}\frac{1}{k}\sum_{t=0}^{k-1}f(\mathbf{w}_{k})\Big{]}\leq\frac{7}{\sqrt[4]{k}}\min_{\mathbf{w}\in\mathcal{C}_{\star}}\left\|{\mathbf{w}_{0}-\mathbf{w}}\right\|^{2}\,.$ ($\tau$ without-replacement)
Similarly, by Lemma 33,
$\mathbb{E}_{\tau}f_{\tau(k+1)}(\mathbf{w}_{k})\triangleq\mathbb{E}_{\tau}\tfrac{1}{2}\left\|\mathbf{w}_{k}-{\Pi}_{\tau\left(k+1\right)}\left(\mathbf{w}_{k}\right)\right\|^{2}\leq\frac{e}{2\sqrt[4]{k}}\min_{\mathbf{w}\in\mathcal{C}_{\star}}\left\|{\mathbf{w}_{0}-\mathbf{w}}\right\|^{2}\,.$ ($\tau$ without-replacement)
Combining the last two displays with Proposition 17, we now obtain:
$\mathbb{E}_{\tau}\bar{f}(\mathbf{w}_{k})$ $\triangleq\mathbb{E}_{\tau}\Big{[}\frac{1}{2T}\sum_{m=1}^{T}\left\|{\mathbf{w}_{k}-{\Pi}_{m}(\mathbf{w}_{k})}\right\|^{2}\Big{]}$ ($\tau$ without-replacement)
$=\frac{k}{T}\mathbb{E}_{\tau}\hat{f}_{0:k-1}(\mathbf{w}_{k})+\frac{T-k}{2T}\mathbb{E}_{\tau}\left\|\mathbf{w}_{k}-{\Pi}_{\tau\left(k+1\right)}\left(\mathbf{w}_{k}\right)\right\|^{2}$
[…]
which proves the without-replacement case and thus completes the proof.
  • Recall Theorem 15.Under a random ordering, with or without replacement, over $T$ jointly separable tasks, the expected forgetting of the weakly-regularized Scheme 5 (at $\lambda\to 0$) after $k\geq 1$ iterations (without replacement: $k\in\left[{T}\right]$) is bounded as
    $\mathbb{E}_{\tau}\big{[}F_{\tau}({k})\big{]}\leq\frac{7\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{\sqrt[4]{k}}\,,\quad\text{where }\mathbf{w}_{\star}\triangleq{\min}_{\mathbf{w}\in\mathcal{C}_{1}\cap\dots\cap\mathcal{C}_{T}}\left\|{\mathbf{w}_{0}-\mathbf{w}}\right\|^{2}\,.$
    Proof.
    We adopt the same notation as used above:
    $f_{m}(\mathbf{w})$ $\triangleq\frac{1}{2}\left\|{\mathbf{w}-{\Pi}_{m}(\mathbf{w}))}\right\|^{2}$
    $\bar{f}(\mathbf{w})$ $\triangleq\mathbb{E}_{m\sim\textrm{Unif}([T])}f_{m}(\mathbf{w})=\frac{1}{2T}\sum_{m=1}^{T}\left\|{\mathbf{w}-{\Pi}_{m}(\mathbf{w})}\right\|^{2}\,.$
    For $\tau$ sampled with replacement, by Lemma 38 (given below) and the with-replacement result (inside the proof) of Theorem 14, we have
[…]
    For $\tau$ sampled without replacement, as argued in Theorem 14, by Lemma 33:
    $\mathbb{E}_{\tau}f_{\tau(k+1)}(\mathbf{w}_{k})\leq\frac{\frac{e}{2}\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{\sqrt[4]{k}}\,,$
    and thus by Lemma 34,
    $\mathbb{E}_{\tau}[F_{\tau}(k)]=\mathbb{E}\hat{f}_{0:k-1}(\mathbf{w}_{k})\leq\left({\frac{e}{\sqrt[4]{k}}+\frac{4}{k}}\right)\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}\leq\frac{7\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{\sqrt[4]{k}}\,.$
    which completes the proof.
Lemma 38.
Consider with-replacement SGD Eq. 2 with step size $\eta\leq 2/\beta$, and define, for every $0\leq T$, $\hat{f}_{0:T}(\mathbf{w})\triangleq\frac{1}{T+1}\sum_{t=0}^{T}f(\mathbf{w};i_{t})$. For all $1\leq T$, the following holds:
$\mathbb{E}\hat{f}_{0:T-1}(\mathbf{w}_{T})$ $\leq 2\mathbb{E}\bar{f}(\mathbf{w}_{T})+\frac{4\beta^{2}\eta\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{T}\,.$
Proof.
Our proof here mostly follows the proof of Lemma 34. Recall that from Eq. (8), any $\beta$-smooth realizable function $h:\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ holds that $\left|h(\tilde{\mathbf{w}})-h(\mathbf{w})\right|\leq h(\mathbf{w})+\beta\left\|{\tilde{\mathbf{w}}-\mathbf{w}}\right\|^{2}$. Denote $f_{t}\triangleq f(\cdot;i_{t})$ for all $t\in\left\{0,.\hskip 0.8pt\!.\hskip 0.8pt\!.,T\right\}$. Now, by the standard stability $\iff$ generalization argument (Shalev-Shwartz et al., 2010; Hardt et al., 2016), and denoting by $\mathbf{w}_{\tau}^{(i)}$ the SGD iterate after $\tau$ steps on the training set where the $i$th example was resampled as $j_{i}$:
$\left|\mathbb{E}\left[\bar{f}(\mathbf{w}_{T})-\hat{f}_{0:T-1}(\mathbf{w}_{T})\right]\right|$ $=\Big{|}{\frac{1}{T}\sum_{i=0}^{T-1}\mathbb{E}_{j_{i}\sim\mathcal{D}}\left[f(\mathbf{w}_{T};j_{i})-f(\mathbf{w}_{T}^{(i)};j_{i})\right]}\Big{|}$
[…]
Next, we bound $\left\|{\mathbf{w}_{T}^{(i)}-\mathbf{w}_{T}}\right\|^{2}$. By the non-expansiveness of gradient steps in the convex and $\beta$-smooth regime when $\eta\leq 2/\beta$ (see Lemma 3.6 in Hardt et al., 2016):
$\tau\leq i$ $\implies\left\|{\mathbf{w}_{\tau}^{(i)}-\mathbf{w}_{\tau}}\right\|=0,$
$i<\tau$ $\implies\left\|{\mathbf{w}_{\tau+1}^{(i)}-\mathbf{w}_{\tau+1}}\right\|^{2}\leq\left\|{\mathbf{w}_{i+1}^{(i)}-\mathbf{w}_{i+1}}\right\|^{2}.$
Further,
$\left\|{\mathbf{w}_{i+1}^{(i)}-\mathbf{w}_{i+1}}\right\|^{2}$ $=\left\|{\mathbf{w}_{i}^{(i)}-\eta\nabla f_{j_{i}}(\mathbf{w}_{i}^{(i)})-\left({\mathbf{w}_{i}-\eta\nabla f_{i}(\mathbf{w}_{i})}\right)}\right\|^{2}$
$\left[\begin{subarray}{c}\mathbf{w}_{i}^{(i)}=\mathbf{w}_{i}\end{subarray}\right]$ $=\eta^{2}\left\|{\nabla f_{j_{i}}(\mathbf{w}_{i}^{(i)})-\nabla f_{i}(\mathbf{w}_{i})}\right\|^{2}$
$\left[\begin{subarray}{c}\text{Jensen}\end{subarray}\right]$ $\leq 2\eta^{2}\big{\|}{\nabla f_{j_{i}}(\mathbf{w}_{i}^{(i)})}\big{\|}^{2}+2\eta^{2}\big{\|}{\nabla f_{i}(\mathbf{w}_{i})}\big{\|}^{2}$
$\left[\begin{subarray}{c}\text{smoothness,}\\ \text{non-negativity}\end{subarray}\right]$ $\leq 4\beta\eta^{2}f_{j_{i}}(\mathbf{w}_{i}^{(i)})+4\beta\eta^{2}f_{i}(\mathbf{w}_{i})\,.$
Therefore,
$\mathbb{E}\left\|{\mathbf{w}_{T}^{(i)}-\mathbf{w}_{T}}\right\|^{2}\leq\mathbb{E}\left\|{\mathbf{w}_{i+1}^{(i)}-\mathbf{w}_{i+1}}\right\|^{2}$ $\leq 4\beta\eta^{2}\mathbb{E}f_{j_{i}}(\mathbf{w}_{i}^{(i)})+4\beta\eta^{2}\mathbb{E}f_{i}(\mathbf{w}_{i})=8\beta\eta^{2}\mathbb{E}f_{i}(\mathbf{w}_{i})\,.$
Now,
$\frac{\beta}{T}\sum_{i=0}^{T-1}\mathbb{E}\left\|{\mathbf{w}_{T}^{(i)}-\mathbf{w}_{T}}\right\|^{2}$ $\leq{12\beta^{2}\eta^{2}}\,\mathbb{E}\left[\frac{1}{T}\sum_{i=0}^{T-1}f_{i}(\mathbf{w}_{i})\right].$
Summarizing, we have shown that:
$\left|\mathbb{E}\left[\bar{f}(\mathbf{w}_{T})-\hat{f}_{0:T-1}(\mathbf{w}_{T})\right]\right|$ $\leq\mathbb{E}\bar{f}(\mathbf{w}_{T})+\frac{\beta}{T}\sum_{i=0}^{T-1}\mathbb{E}\left\|{\mathbf{w}_{T}^{(i)}-\mathbf{w}_{T}}\right\|^{2}$
$\leq\mathbb{E}\bar{f}(\mathbf{w}_{T})+{8\beta^{2}\eta^{2}}\,\mathbb{E}\left[\frac{1}{T}\sum_{i=0}^{T-1}f_{i}(\mathbf{w}_{i})\right].$
Finally, by the regret bound given in Lemma 11, i.e., $\sum_{i=0}^{T-1}f_{i}(\mathbf{w}_{i})\leq\frac{\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{2\eta(2-\eta\beta)}$, we have
$\left|\mathbb{E}\left[\bar{f}(\mathbf{w}_{T})-\hat{f}_{0:T-1}(\mathbf{w}_{T})\right]\right|$ $\leq\mathbb{E}\bar{f}(\mathbf{w}_{T})+\frac{4\beta^{2}\eta\left\|{\mathbf{w}_{0}-\mathbf{w}_{\star}}\right\|^{2}}{(2-\eta\beta)T}\,.$
and the result follows.